### v1

# A Survey on Web Application Testing: A Decade of Evolution

Abstract.
As one of the most popular software applications, a web application is a program, accessible through the web, to dynamically generate content based on user interactions or contextual data, for example, online shopping platforms, social networking sites, and financial services. Web applications operate in diverse environments and leverage web technologies such as HTML, CSS, JavaScript, and Ajax, often incorporating features like asynchronous operations to enhance user experience. Due to the increasing user and popularity of web applications, approaches to their quality have become increasingly important. Web Application Testing (WAT) plays a vital role in ensuring web applications’ functionality, security, and reliability. Given the speed with which web technologies are evolving, WAT is especially important. Over the last decade, various WAT approaches have been developed. The diversity of approaches reflects the many aspects of web applications, such as dynamic content, asynchronous operations, and diverse user environments. This paper provides a comprehensive overview of the main achievements during the past decade: It examines the main steps involved in WAT, including test-case generation and execution, and evaluation and assessment. The currently available tools for WAT are also examined. The paper also discusses some open research challenges and potential future WAT work.
Software testing, web application testing, survey
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software verification and validationccs: Information systems Web applicationsccs: General and reference Surveys and overviews
1. Introduction
The Software Development Life Cycle (SDLC) (Shafiq et al., 2021) is composed of several essential phases, including requirements specification, architecture design, coding, testing, and maintenance (Hanna and Ahmad, 2022). Each phase involves specific processes and activities that contribute to the overall development of the software. Among these phases, testing is particularly critical, often accounting for a significant proportion of time and cost for developing a software application. Recent studies (López-Martín, 2022; Bindu Bhargavi and Suma, 2022) indicate that the testing process can consume a range of 10% to 80% of total development time, depending on the complexity of the project. The primary objective of testing is to identify errors that manifest during program execution. An effective testing methodology improves the probability of detecting such errors, for example, by systematically running and examining a wide range of execution paths, including edge cases and fault-prone areas (Li et al., 2014b; Mahdieh et al., 2022).
Since the adoption of the World Wide Web in the early 1990s (Berners-Lee, 1992), web applications have become integral to various sectors, including finance, healthcare, education, and commerce (Doğan et al., 2014). These applications have introduced innovative features and functionalities that facilitate consistent and efficient access to information across diverse sources, significantly transforming how society interacts with digital content (Jazayeri, 2007). Meanwhile, web applications rely on a set of core technologies, including HTML to define the structure of web pages, CSS to style, JavaScript to enable interactivity, and Ajax for asynchronous data exchanges between the client and the server. These technologies are the foundation of modern web applications and are critical in supporting their dynamic and distributed nature (Garrett et al., 2005; Arora and Sinha, 2013; Sampath and Sprenkle, 2016).
As web applications grow increasingly complex and technologies continue to evolve, the requirement for robust and effective testing methodologies has become more crucial atkinson2002towards: It is necessary to continuously evolve and advance testing strategies to ensure the reliability, security, and overall quality of web applications (Li et al., 2014b). Web Application Testing (WAT) (Doğan et al., 2014) has become an essential process in the SDLC of web applications, which is used to evaluate and ensure the quality of Applications Under Test (AUTs) (Karimoddini et al., 2022), making it a cornerstone of modern software quality assurance. WAT aims to ensure that web applications function as intended, meet specified requirements, and are free from vulnerabilities. The scope of WAT goes beyond bug detection and fixing, and includes enhancing user experience, ensuring data security, and maintaining the overall reliability of the web systems. As web technologies evolve, testing methodologies have advanced to address web applications’ increasing complexities and demands.
Web applications have had an important impact on society (Garousi et al., 2013), reshaping various aspects of daily life and industry (Onukrane et al., 2023). Despite their significance, however, WAT reviews have tended to be narrowly focused (Prazina et al., 2023), addressing specific aspects without providing a comprehensive overview (Kousar et al., 2023). One WAT survey paper, by Prazina et al. (Prazina et al., 2023), only concentrates on the testing and analysis methods for automated web layouts. In contrast, significant attention has been devoted to the security aspects of web applications (Rahman and Izurieta, 2022; Aydos et al., 2022; Zhang et al., 2021; Altayaran and Elmedany, 2021; Cui et al., 2020; Wang et al., 2020; Muslihi and Alghazzawi, 2020; Marashdih et al., 2019; Kumar and Sheth, 2016; Rafique et al., 2015; Gupta et al., 2014). Additionally, Gupta et al. (Gupta et al., 2018) focused on the automated generation of regression test cases for web applications.
Early reviews of WAT, such as those by Li et al. (Li et al., 2014b) and Dougan et al. (Doğan et al., 2014), established foundational insights into testing dynamic web environments, focusing on early challenges and techniques. More recently, Balsam and Mishra (Balsam and Mishra, 2024) have provided a review to discuss some challenges and opportunities in WAT. However, this paper focused on only 72 primary studies, which means that many significant contributions and broader advancements, over the past decade, have been overlooked. To address this gap, we aim to provide a comprehensive examination of the latest progress in WAT, detailing core processes, evaluating current tools, and exploring emerging research directions to offer a thorough overview of the field’s evolution over the past decade.
This paper provides a comprehensive examination and discussion of WAT across its entire life cycle. We explore the research status and WAT advances from the perspectives of test case generation and execution, failure diagnosis, evaluation and assessment, regression testing, and available tools. We also examine some key open research challenges and discuss potential directions for future WAT work.
The rest of this paper is organized as follows: Section 2 provides background information on the formal definition of WAT and the fundamental testing techniques employed in the field. Section 3 outlines the survey methodology, including presenting the six research questions, the literature retrieval process, and the statistical analysis of the retrieved studies. Sections 4 to 9 address each of the research questions (RQ1 to RQ6), respectively. Finally, Section 10 concludes the paper by summarizing key findings and proposing directions for future research.
2. Background
This section presents some preliminary WAT concepts and provides an overview of web applications and WAT, covering some core definitions and relevant testing techniques.
2.1. Web Applications
A web application (Jazayeri, 2007) is a software system that facilitates user interaction through a combination of front-end and back-end components, over a network, typically accessed through a web browser. The front-end, or client-side, renders the user interface and handles user inputs. It consists of web pages constructed using technologies such as HTML, CSS, and JavaScript. This layer is designed to provide a responsive and interactive user experience, enabling seamless communication between the user and the application. The back-end, or server-side, encompasses the core functionalities of the application: These include data processing, business logic execution, and database management. The back-end is the key to the application: It processes requests from the front-end and ensures that the correct data and functionalities are provided to the user. Communication between the front-end and back-end typically relies on protocols such as HTTP and HTTPS, which support smooth data exchange and maintain the application’s state (Kumar et al., 2017; Neumann et al., 2018). Together, these components provide an integrated system that supports a wide range of user interactions, from simple content browsing to complex transaction processing.
Web applications can be either static or dynamic. Static web applications have fixed content that remains unchanged regardless of user interactions, making them suitable for delivering straightforward information, such as in digital brochures or basic company websites (Vaidyanathan and Mautone, 2009). Dynamic web applications, in contrast, provide a more interactive experience, with content that adapts to user inputs, interactions, or real-time data (Sampath and Sprenkle, 2016). The dynamic nature of these applications allows for greater flexibility and user engagement, making them well-suited for complex platforms like social media networks, e-commerce sites, and online service portals that require continuous interaction and real-time responsiveness (Fulcini and Ardito, 2022).
2.2. Web Application Testing
Web applications have become a fundamental part of our daily lives, making thorough testing a vital process to ensure their functionality and dependability. Effective testing ensures that these applications not only meet the required functionality but also have sufficient security, and good performance and reliability. Thorough testing of web applications can help to identify and address potential vulnerabilities, performance bottlenecks, and functional issues: This enhances the user experience and helps to safeguard sensitive data, which is vital for maintaining trust in web-based services. Comprehensive testing ultimately contributes to the development of robust, dependable web applications that can adapt to the dynamic demands of users and the evolving technological landscape.
WAT involves several key steps:
  • Analysis of the functional and non-functional requirements of the web application to guide the development of comprehensive test cases.
  • Generation of test data according to the testing methodology (such as model-based or security testing), and appropriate design of test cases, to ensure a broad coverage of potential scenarios.
  • Execution of the test cases by sending HTTP requests to the web application, and comparison of the responses with the expected outcomes to identify inconsistencies.
  • Documentation of the test results, analysis of any issues, and application of regression testing to ensure that recent changes have not introduced new defects.
  • Implementation of continuous testing and monitoring to maintain the application’s stability and security as it evolves.
3. Methodology
In this section, we present the six research questions (RQs) related to WAT that structured our study, and this paper. Our WAT survey involved a systematic and structured methodology, guided by the published frameworks of Huang et al. (Huang et al., 2021) and Webster et al. (Webster and Watson, 2002). Our findings’ presentation draws from recent surveys on related topics, including the testing of RESTful APIs (Golmohammadi et al., 2023), mutation analysis (Jia and Harman, 2010), and metamorphic testing (Segura et al., 2016; Chen et al., 2018). The following sections detail key aspects of the review process and results.
3.1. Research Questions
The goal of this survey paper was to identify and categorize the available WAT information. To achieve this, the following six RQs guided the study:
  • RQ1: How has WAT evolved, and what WAT topics have been examined in published studies?
  • RQ2: What methods and approaches have been used for WAT test case generation?
  • RQ3: How are test cases executed in WAT?
  • RQ4: What metrics are used to evaluate WAT? How are they used?
  • RQ5: What tools are available to support WAT?
  • RQ6: What are the challenges and future work related to WAT?
An answer to RQ1 should provide an understanding of the WAT literature and the distribution of its key topics. RQ2 prompted an examination of the various methods and approaches used for WAT test case generation, revealing the application of a diversity of techniques. RQ3 focuses on the execution of WAT test cases, aiming to clarify the processes and practices involved. RQ4 should reveal the metrics and evaluation frameworks used in WAT, enabling an understanding of how the WAT effectiveness and efficiency are assessed. RQ5 should reveal the currently available WAT tools, offering a comprehensive overview of the technological resources for WAT. Finally, based on the insights gathered from the other RQs, RQ6 shall identify unresolved challenges and propose potential directions for future WAT research. Figure 1 presents an overview of the structure of this paper.
Figure 1. Structure of this survey paper.
3.2. Literature Search and Selection
Similar to previous survey studies (Huang et al., 2021; Webster and Watson, 2002; Golmohammadi et al., 2023; Harman et al., 2015; Anand et al., 2013), we selected the following online repositories of technical research literature:
  • ACM Digital Library (ACM)
  • Elsevier Science Direct (Elsevier)
  • IEEE Xplore Digital Library (IEEE)
  • Springer Online Library (Springer)
  • Wiley Online Library (Wiley)
These repositories were chosen for their extensive collections of conference, symposium, and workshop papers, as well as for their access to multiple journals that are considered highly relevant to WAT (Huang et al., 2021). Our survey examined computer science papers published between January 1, 2014, and December 31, 2023.
Once the literature repositories were identified, search strategies were developed and tailored for each of them, using WAT-specific terminology and formatting. Given the unique characteristics and limitations of the advanced search functions in each repository (such as Elsevier’s restriction on using a maximum of eight Boolean connectors per field), the search methods were designed to strictly adhere to the repository rules and constraints. To ensure comprehensive coverage of all relevant literature, multi-dimensional and multi-layered filtering keywords were used to maximize the retrieval of the WAT materials. Table 1 illustrates the specific search queries for five selected digital libraries.
Table 1. Selected Digital Libraries, with Search Queries
No. Name Search Query
1 ACM [[Title: ‘‘web’’] OR [Title: ‘‘browser based’’]] AND [[Title: ‘‘test’’] OR [Title: ‘‘testing’’] OR [Title: ‘‘measure’’] OR [Title: ‘‘measurement’’] OR [Title: ‘‘measuring’’] OR [Title: ‘‘check’’] OR [Title: ‘‘checking’’] OR [Title: ‘‘detect’’] OR [Title: ‘‘detecting’’] OR [Title: ‘‘detection’’]]
2 Elsevier ((‘‘web’’) OR (‘‘browser based’’)) AND ((‘‘test’’) OR (‘‘testing’’) OR (‘‘measure’’) OR (‘‘check’’) OR (‘‘checking’’) OR (‘‘detect’’) OR (‘‘detecting’’))
3 IEEE (‘‘Document Title’’:‘‘web’’ OR ‘‘Document Title’’:‘‘browser based’’) AND (‘‘Document Title’’:‘‘test’’ OR ‘‘Document Title’’:‘‘testing’’ OR ‘‘Document Title’’:‘‘measure’’ OR ‘‘Document Title’’:‘‘measurement’’ OR ‘‘Document Title’’:‘‘measuring’’ OR ‘‘Document Title’’:‘‘check’’ OR ‘‘Document Title’’:‘‘checking’’ OR ‘‘Document Title’’:‘‘detect’’ OR ‘‘Document Title’’:‘‘detecting’’ OR ‘‘Document Title’’:‘‘detection’’)
4 Springer ‘‘web’’ AND (‘‘test’’ OR ‘‘testing’’ OR ‘‘measure’’ OR ‘‘measurement’’ OR ‘‘measuring’’ OR ‘‘check’’ OR ‘‘checking’’ OR ‘‘detect’’ OR ‘‘detecting’’ OR ‘‘detection’’)
5 Wiley ‘‘(‘‘web’’) OR (‘‘browser based’’)’’ in Title and ‘‘(‘‘test’’) OR (‘‘testing’’) OR (‘‘measure’’) OR (‘‘measurement’’) OR (‘‘measuring’’) OR (‘‘check’’) OR (‘‘checking’’) OR (‘‘detect’’) OR (‘‘detecting’’) OR (‘‘detection’’)’’ in Title
The survey included studies written in English that were focused on WAT, but excluded academic theses (e.g., master’s and doctoral dissertations), books, posters, and previous survey studies. In addition, some processes relevant to WAT, such as failure diagnosis and regression testing for web applications, were excluded from the survey. The inclusion criteria were:
  • The paper was written in English.
  • The paper was related to WAT.
  • The paper was not a thesis.
  • The paper was not a survey or a systematic literature review.
  • The paper was freely available in open access.
After duplicates were removed and the exclusion criteria applied, the initial set of candidate studies was significantly refined. A snowballing approach (Golmohammadi et al., 2023), which involved examining the references lists of in the selected studies, led to the identification and inclusion of several additional papers. Table 2 summarized the details of the search and filtering process. Eventually, 314 papers were included for the preliminary review and statistical analysis.
Although some papers may have been omitted due to the focus on a subset of reputable publishers, we are confident that the overall trends we report on are accurate and provide a fair representation of the current state of the art in WAT.
3.3. Data Extraction and Collection
Key information was gathered from each study regarding the research motivation, contributions to the field, empirical-evaluation details, common misconceptions, and remaining challenges in WAT. These were reviewed and verified by co-authors of the survey, minimizing the risk of overlooking critical information, and reducing potential errors in the analyses.
We organized the RQ data collection methods as follows:
  • RQ1: Fundamental information for each paper such as publication year, type of paper, and paper source.
  • RQ2: Testing AUT, testing objective, and test case generation techniques.
  • RQ3: Execution methods, tools and platforms, environments, tested items, dataset URL, and open-source status.
  • RQ4: Evaluation metrics, methods and approaches, and tools and techniques.
  • RQ5: Tool name, type, main features and capabilities, URL, and open-source status.
  • RQ6: Overview of the remaining challenges.
Table 2. Selection Results of Primary Studies
No. Name
No. of studies from the search
keywords-based results
No. of studies excluded
based on the selection criteria
No. of studies after
the selection criteria
1 ACM 232 138 94
2 Elsevier 27 19 8
3 IEEE 343 171 172
4 Springer 65 33 32
5 Wiley 19 11 8
Total 686 372 314
4. ANSWER TO RQ1: Publication Evolution and Distribution
This section addresses RQ1 through a detailed analysis of publication trends and their distribution across the literature sources. Our methodology provides a comprehensive analysis of the evolution and distribution of WAT research over the past decade.
4.1. Publication Trends
We collected publication year data from 314 papers (as illustrated in Figure LABEL:Fig:publication_trends) to highlight trends in WAT research from 2014 to 2023.
Figure LABEL:Fig:number_of_publications_per_year shows the annual number of publications, and Figure LABEL:Fig:cumulative_number_of_publications_per_year shows the cumulative total. Figure LABEL:Fig:number_of_publications_per_year shows the fluctuating (but relatively steady) nature of WAT-related publications from 2014 to 2023. The annual publication counts exhibit some variation over the decade, with 32 publications in 2014, followed by slight fluctuations: 35 in 2015, 25 in 2016, 27 in 2017, 29 in 2018, and 34 in 2019. In 2020, the count was 31, which increased to 38 in 2021. This was followed by a decrease to 28 in 2022 and a subsequent rise to 35 in 2023. This pattern reflects a generally stable trend in WAT research output over the years.
The cumulative publication data presented in Figure LABEL:Fig:cumulative_number_of_publications_per_year reflects a consistent growth in research contributions. The trend in Figure LABEL:Fig:cumulative_number_of_publications_per_year suggests that WAT has maintained a steady pace of development, driven by ongoing technological advancements and sustained research interest in the field.
4.2. Paper Source and Publisher
The papers analyzed in this study were from various journals and conferences. As shown in, the majority of the publications were from conferences (71%), with a smaller proportion being from journals (29%). The data presented in Figure LABEL:Fig:venues_distribution_per_year reveals distinct patterns in publication volumes over the years. For instance, conference papers exhibited fluctuations, starting at 25 in 2014, rising slightly to 27 in 2015, and then decreasing to 17 in 2018 before recovering to 26 in 2023. In contrast, journal publications, while also fluctuating, reached their peak of 15 papers in 2019, demonstrating a steadier but less pronounced growth compared to conference papers.
5. ANSWER TO RQ2: Test Case Generation
This section addresses RQ2 by systematically examining key aspects of WAT test case generation, including the Applications Under Test (AUTs) and datasets, testing objectives, and test case generation methods.
5.1. AUTs and Datasets
5.1.1. AUTs
This section reviews the types of AUTs used in WAT, referencing the selected research papers. In WAT, AUTs refer to the web platforms or systems being evaluated (for functionality, performance, security, and usability). The diversity of WAT AUTs emphasizes the need for tailored testing approaches to address specific challenges across different domains: E-commerce platforms, for example, require stringent security measures, whereas government portals may focus on accessibility.
  • E-commerce Platforms: E-commerce applications such as Magento, Shopify, and custom-built online stores are popular WAT AUTs. This may be due to their complex transactional workflows, which involve user authentication, shopping carts, and payment gateways (Pratama and Rhusuli, 2022). Testing often focuses on functional correctness and security, with particular attention to vulnerabilities to potential attacks like SQL Injection (SQLi) and Cross-Site Scripting (XSS) (Bozic et al., 2015a). Performance testing is also essential to ensure that the platforms can handle high traffic volumes during peak times (Wan et al., 2019).
  • Content Management Systems (CMSs): CMS platforms such as WordPress, Drupal, and Joomla are frequently chosen as AUTs because of their modular and extensible nature, which can potentially make them vulnerable, through third-party plugins (Garn et al., 2014; Arcuri et al., 2023). Testing typically emphasizes plugin vulnerabilities, core stability, and validating performance after updates (Li et al., 2014c).
  • Government and Public Service Portals: Government web applications, such as university and public service portals, are often tested for accessibility, usability, and security (Offutt and Thummala, 2019). These platforms often need to comply with accessibility standards such as the Web Content Accessibility Guidelines (WCAG) (Ara et al., 2024), making accessibility testing critical (Vithanage and Jeyamohan, 2016). Security testing emphasizes the prevention of unauthorized access and data breaches, as illustrated by frameworks targeting GDPR compliance in web portals (Amanzholova et al., 2021).
  • Social Media Platforms: Social media platforms, like Facebook and Twitter, pose unique challenges for WAT, due to their real-time interaction and dynamic content. Testing for these applications often examines performance under heavy user loads, cross-browser compatibility, and session management vulnerabilities (Fan et al., 2023; Appelt et al., 2018).
  • Web Application Firewalls (WAFs): WAFs play a key role in defending web applications from attacks like SQLi, XSS, and Cross-Site Request Forgery (CSRF) (Pelivani and Cico, 2021; Zhang et al., 2015a). Testing WAFs involves simulating real-world attack scenarios to evaluate how effectively they distinguish between legitimate and malicious traffic (Zhao and Gong, 2015; Chang et al., 2023).
Table 3. An overview of datasets.
Dataset Name Overview Resource Link Reference
Software Assurance Reference Dataset (SARD) It is a dataset consisting of PHP test cases, where each test case represents a vulnerable or non-vulnerable SQL-related code snippet. The test cases include examples of user input propagating through the application code until a sensitive function (e.g., a database query) is executed. https://samate.nist.gov/SARD/index.php (Fidalgo et al., 2020)
Near-Duplicate Study DataSet It is a dataset of about 100,000 annotated pairs of web pages from open-source web applications and real websites. Each pair is labeled as distinct, near-duplicate, or clone to facilitate the evaluation of near-duplicate detection methods. https://zenodo.org/records/3376730  (Corazza et al., 2021)
Website Screenshots Dataset from Roboflow It is a dataset of 1,206 screenshots from popular websites with 54,215 UI element annotations such as buttons, text, images, links, titles, fields, iframes, and labels. This dataset is used to train object detection models and evaluate the visual consistency of web applications. https://public.roboflow.com/object-detection/website-screenshots  (Aditya et al., 2023)
PhishTank dataset It is a widely used dataset containing known phishing URLs. The dataset provides examples of real phishing attempts that the system uses to test its effectiveness. http://www.phishtank.com  (Nirmal et al., 2021)
Web injection publicly available datasets The datasets consist of various HTTP request samples labeled according to the type of attack they represent, including SQL injection, XSS, command injection, and path traversal. SQL Injection: (1) https://www.kaggle.com/syedsaqlainhussain/sql-injection-dataset,
(2) https://github.com/mhamouei/rat_datasets, XSS Injection: https://www.kaggle.com/syedsaqlainhussain/cross-site-scripting-xss-dataset-for-deep-learning, HTTP Parameters: https://github.com/Morzeux/HttpParamsDataset
 (Román-Gallego et al., 2023; Amouei et al., 2021)
Mitch Framework HTTP Request Dataset It is a dataset based on the usage of the Mitch framework, which contains 5,828 HTTP requests but is reduced to 1901 records after balancing. https://github.com/alviser/mitch  (Beltrano et al., 2023)
NASA 1995 Apache Web Log It is a dataset containing log files from NASA web servers, recording HTTP requests from real users during July 1995. https://ita.ee.lbl.gov/html/contrib/NASA-HTTP.html  (Pavanetto and Brambilla, 2020)
DBpedia It is a structured knowledge base that extracts information from Wikipedia. It contains millions of entities represented as RDF triples. https://www.dbpedia.org/  (Mariani et al., 2014)
fuzzdb It is an open-source database containing a large number of attack patterns. https://github.com/fuzzdb-project/fuzzdb  (Zhang et al., 2017)
ATO-data It is a collection of keystroke dynamics data collected under controlled conditions in a lab and anonymized data from 187,332 sessions of financial web pages, which supports analyzing the changes in different user behaviors and is used to evaluate the effectiveness of imposter detection models. https://github.com/mluckner/ATO-data.git  (Grzenda et al., 2023)
Web Accessibility Evaluation Data It is an automated web accessibility assessment of nearly 3 million web pages obtained between March and September 2021 using Qual Web. https://zenodo.org/records/7494722  (Martins and Duarte, 2023)
PHP-Webshell-Dataset It is a dataset containing 2,917 samples from 17 webshell collection projects. https://github.com/Cyc1e183/PHP-Webshell-Dataset  (Gogoi et al., 2022)
PHP Vulnerability Test Suite It is a dataset containing 9408 PHP source code samples, including 5600 safe samples and 3808 unsafe samples, which is widely used for the detection of XSS vulnerabilities and the training of prediction models. https://github.com/stivalet/PHP-Vulnerability-test-suite  (Gupta et al., 2015)
Datasets of three common web attack payloads (SQLi, XSS, RCE) It is a dataset generated by attack grammar, including SQL injection, cross-site scripting, and remote command execution attack payloads, used to train and test the performance of GPTFuzzer on WAF. https://github.com/hongliangliang/gptfuzzer  (Liang et al., 2023)
Joomla! and MantisBT test suite It is a data set that contains multiple representative use cases, covering functions such as article management and user management. It has 47 manually written test cases and 453 test steps. https://zenodo.org/record/4973219  (Kirinuki et al., 2021)
GHTraffic It is a dataset of significant size comprising HTTP transactions extracted from GitHub data (i.e., from 04 August 2015 GHTorrent issues snapshot) and augmented with synthetic transaction data. https://zenodo.org/record/1034573  (Bhagya et al., 2019)
5.1.2. Datasets
Datasets are essential in WAT, providing the inputs needed to simulate real-world usage and evaluate application behavior across key areas such as security, functionality, and performance. The datasets are often comprised of user data, predefined test cases, or automatically generated data. They support a range of testing scenarios, including edge cases, aiming to identify potential vulnerabilities and performance bottlenecks.
Each WAT dataset was designed to meet specific testing objectives, allowing for targeted assessments of various aspects. For example, the Software Assurance Reference Dataset (SARD) (Fidalgo et al., 2020) can help to identify security vulnerabilities, such as SQLi, thus contributing to more secure software development practices. The PhishTank dataset (Nirmal et al., 2021), which contains real-world phishing URLs, can support evaluation of phishing detection systems, enhancing web application security.
Datasets like the Website Screenshots Dataset (Aditya et al., 2023) facilitate the assessment of visual consistency and UI accuracy, helping to ensure an optimal user experience. The NASA web log data (Pavanetto and Brambilla, 2020) provides real-world HTTP requests, offering valuable insights into traffic patterns and server performance under various loads. DBpedia (Mariani et al., 2014) supports the validation of semantic web applications: It provides structured data that can be used to help verify data integration and usage.
These datasets enable testers to adopt focused testing strategies that ensure a thorough evaluation of web applications across multiple dimensions. Table 3 summarizes key datasets used in WAT, each with a distinct focus, ranging from security testing.
5.2. Test Objectives
WAT test objectives include detecting bugs, ensuring security, evaluating performance, assessing the functionality of Graphical User Interfaces (GUIs), and identifying vulnerabilities (such as for XSS and SQLi).
Figure 3. Proportions of test objectives in surveyed papers.
Figure 3 shows the proportion of the different objectives across the surveyed literature. Web vulnerability detection, including XSS and SQL Injection detection, accounted for the largest proportion (29.94%), emphasizing the necessity of addressing critical security risks. Security testing represented 27.71%, highlighting its importance in protecting applications from security threats. GUI testing accounted for 19.43%. Performance evaluation constituted 12.42%, and bug/defect detection represented 10.51%. The following sections provide a detailed analysis of each testing objective.
5.2.1. Bug/Defect Detection
Detecting and correcting defects is essential for maintaining web application integrity. Techniques such as combining unsupervised learning with embedding layers effectively optimize test case prioritization while reducing redundancy (Demirel and Sandikkaya, 2023). Modern web applications often involve intricate interdependencies between components, requiring advanced techniques to detect defects across diverse states and modules. This complexity necessitates strategies such as dynamic test case generation and static analysis to ensure robust coverage (Maskur and Asnar, 2019).
Static code analysis can complement regression testing, and may identify issues early in development. Tools that use static analysis (such as taint analysis) detect potential errors in the source code without execution (Maskur and Asnar, 2019; Nunes et al., 2019). For example, the Q-learning-based Selenium-Java tool dynamically explores web GUI elements and learns optimal test paths, significantly improving testing coverage for complex systems (Fan et al., 2023).
5.2.2. Security Testing
The goal of security testing is to ensure that web applications remain protected against unauthorized access, data breaches, and other security threats. Security testing provides a systematic framework for evaluating the robustness of applications through the identification and mitigation of vulnerabilities before and during runtime (Medeiros et al., 2016).
To achieve comprehensive protection, security testing employs a combination of methodologies. Static analysis examines source code for vulnerabilities, such as unhandled exceptions and unsanitized inputs, enabling early detection of issues during development (Zhao and Gong, 2015). Dynamic analysis evaluates applications during runtime, simulating potential attack scenarios to uncover vulnerabilities that static analysis might overlook (Pellegrino et al., 2015). Hybrid analysis combines static and dynamic methods to address their individual limitations, offering enhanced coverage and reducing false positives (Guan et al., 2021). Automated tools, such as Burp Suite, enhance the efficiency of dynamic analysis by enabling real-time evaluation of application defenses (Garn et al., 2014).
5.2.3. Performance Evaluation
Performance evaluation aims to ensure that an application remains efficient and stable under varying load conditions (Pelivani and Cico, 2021). Essential to this are load testing and stress testing, which simulate diverse user traffic scenarios to help developers identify performance bottlenecks, especially in environments with high concurrency (Pelivani and Cico, 2021; Lounis et al., 2014).
Advanced tools such as JMeter (Kiran et al., 2015; Tiwari et al., 2023) and LoadRunner (Cai and Hu, 2014) can automate measurement of some important metrics, including response times, throughput, and system resource utilization (e.g., CPU, memory, disk I/O) (Fan et al., 2023). They can do this across a spectrum of load levels (Fan et al., 2023). These tools can provide comprehensive insights into system behavior under stress, helping the optimization of the system architecture, and ensuring reliability during peak load conditions (Leotta et al., 2020).
5.2.4. GUI Testing
The aim of GUI testing is to ensure that the application’s user interface remains consistent, in terms of both functionality and user experience, across different devices and browser environments (Zhang et al., 2015b). Automated tools can simulate user interactions, exploring the GUI’s responsiveness under diverse conditions (Lin et al., 2017): Cross-browser testing with tools like Selenium Grid (Kiranagi and Shyam, 2017), for example, can help to ensure compatibility across different platforms (Pelivani and Cico, 2021). Meanwhile, Scout uses gamified features (Fulcini and Ardito, 2022), including progress tracking and dynamic error injection, to encourage systematic exploration and improve test coverage.
Image comparison and computer vision technologies are also extensively used to detect interface changes, and to identify potential UI errors (Leotta et al., 2020). These tools capture and compare screenshots from different stages of testing, ensuring that updates do not introduce unintended changes to the user interface (Zhao and Gong, 2015). The recent integration of AI in testing can further enhance efficiency by learning from user interactions and predicting areas that may be likely to fail (Fan et al., 2023; Chang et al., 2023).
5.2.5. Web Vulnerability Detection
Web vulnerability detection focuses on identifying and mitigating specific threats in web applications, such as Cross-Site Request Forgery (CSRF), directory traversal, Cross-Site Scripting (XSS), and SQL Injection (SQLi). These vulnerabilities, if left unaddressed, can compromise application functionality, data security, and user privacy (Ben Jaballah and Kheir, 2016; Zhang et al., 2015a).
To detect these vulnerabilities, static analysis examines source code to identify input validation flaws and hardcoded credentials, enabling early detection of coding issues (Zhao and Gong, 2015; Lounis et al., 2014). Dynamic analysis simulates attack scenarios during runtime, evaluating the application’s behavior under unexpected interactions (Ben Jaballah and Kheir, 2016). Hybrid analysis combines the strengths of static and dynamic approaches, providing enhanced detection accuracy and reducing false positives (Guan et al., 2021).
Modern advancements in vulnerability detection integrate machine learning and deep learning techniques. Methods based on Recurrent Neural Networks (RNNs) analyze input and output patterns to detect malicious behaviors, demonstrating high effectiveness in identifying XSS and SQLi attacks (Guan et al., 2021). Additionally, OWASP ZAP is widely used to automate the identification of vulnerabilities, offering strong support for detecting SQLi and XSS (Zhang et al., 2015a).
5.3. Generation Methods
WAT can make use of many test case generation methods, including: Model-Based (Dias Neto et al., 2007), Combinatorial and Optimization-Based (Nie and Leung, 2011), Learning-Based (Alagarsamy et al., 2024), Data and Content-Driven (Elbaum et al., 2008), Crawling and Exploration-Based (Mesbah et al., 2012), Template and Page Object-Based (Stocco et al., 2016), and Security Test Case Generation techniques (Felderer et al., 2016). These methods offer a broad selection of strategies that can address diverse testing needs.
Figure 4. Proportions of generation methods in surveyed papers.
Figure 4 presents the proportions of some key WAT test generation methods in the surveyed literature. Model-Based Testing (MBT), which uses structural models (e.g., UML, statecharts) that reflect system behavior, is the most prominent, appearing in 27.50% of the papers. Combinatorial and Optimization-Based Testing (20.00%) enhances coverage through input-variable combinations. Data and Content-Driven Testing (20.00%) make use of input validation across diverse scenarios. Security Testing was found in 15.00% of the papers. Learning-Based Testing (10.00%), Template and Page Object-Based Testing (5.00%), and Crawling and Exploration-Based Testing (2.50%) were less commonly reported.
5.3.1. MBT
Model-Based Testing (MBT) automatically generates test cases, enhancing both efficiency and coverage. UML diagrams, statecharts, ER diagrams, and BPMN models are commonly used to represent the system, and to generate the concrete test cases. Panthi & Mohapatra (Panthi and Mohapatra, 2017) and Suhag & Bhatia (Nabuco and Paiva, 2014) used UML models; Indumathi & Begum (Indumathi and Begum, 2016) used ER diagrams and state transition models for database testing. Akpinar et al. (Akpinar et al., 2020) showed how MBT could enhance coverage of complex interactions by integrating GraphWalker, Selenium, and JUnit. Moura et al. (de Moura et al., 2017) used BPMN models to generate automated test scripts that covered functional test paths in Business Process Management (BPM) applications. Wang et al. (Wang et al., 2016b) proposed an approach based on UML statecharts for generating Web link security test scenarios. Thummala & Offutt (Thummala and Offutt, 2016) used Petri nets to detect concurrency behaviors. Fard et al. (Milani Fard et al., 2014) and Waheed et al. (Waheed et al., 2020) showed how model inference and script automation could enhance the automation and test coverage.
5.3.2. Combinatorial and Optimization-Based Testing
Techniques based on combinatorial or optimization can generate efficient test cases, which can significantly reduce the number of tests while improving coverage. Qi et al. (Qi et al., 2017) introduced an automated approach based on combinatorial strategies: This enhanced the coverage of dynamic pages. Bozic et al. (Bozic et al., 2015b) reported on the effectiveness of IPOG and IPOG-F combinatorial testing algorithms for detecting XSS vulnerabilities. Wang et al. (Wang et al., 2023) and Wang et al. (Wang et al., 2019) applied evolutionary algorithms and combinatorial testing methods to generate test cases covering complex paths in Web applications: Their work demonstrated the potential for genetic algorithms and simulated annealing in Web testing. Wang et al. (Wang et al., 2016a) used combinatorial testing to construct navigation maps for dynamic Web applications, enhancing the test coverage and effectiveness.
5.3.3. AI-based Testing
Testing methods can also use machine learning and artificial intelligence (AI) techniques to analyze and generate Web application behaviors and inputs, automating the creation of effective test cases. Chang et al. (Chang et al., 2023) proposed the WebQT tool based on reinforcement learning, significantly improving testing efficiency and coverage through intelligent path selection. Pavanetto & Brambilla (Pavanetto and Brambilla, 2020) used Generative Adversarial Networks (GANs) to generate highly realistic Web navigation paths, effectively simulating real user operations. Chaleshtari et al. (Chaleshtari et al., 2023) demonstrated the potential of AI techniques in security testing by using mutation testing to detect security vulnerabilities in Web systems. Shahbaz et al. (Shahbaz et al., 2015) introduced a technique for generating test data based on regular expressions and Web searches, producing valid and invalid WAT data. Li et al. (Li et al., 2014a) developed the SymJS framework, which combines symbolic execution and dynamic feedback to improve the coverage and accuracy of JavaScript WAT.
5.3.4. Data and Content-Driven Testing
Data and content-driven testing methods aim to cover all possible input combinations, to comprehensively verify application functionality and security. Hanna & Jaber (Hanna and Jaber, 2019) proposed a method based on analyzing client-side user input fields, effectively covering various input scenarios. Hanna & Munro (Hanna and Munro, 2018) used semantically invalid input data to ensure that Web applications reject inputs that do not meet semantic constraints, thereby enhancing security. Shahbaz et al. (Shahbaz et al., 2015) combined regular expressions with Web searches to automatically generate test cases, verifying string validation logic in Web applications. Nagowah & Kora-Ramiah (Nagowah and Kora-Ramiah, 2017) developed the CRaTCP tool, which automatically extracts control combinations and generates test cases, achieving comprehensive WAT coverage. Mariani et al. (Mariani et al., 2014) utilized the data from applications to automatically generate valid input data (which satisfies the syntactical and semantical requirements) to test complex web applications and ensure the coverage of input combinations.
5.3.5. Crawling and Exploration-Based Testing
Methods based on crawling and exploration are particularly suited for testing the complex navigation and state changes of dynamic Web applications. Fard et al. (Milani Fard et al., 2014) combined automated crawling with manually written test cases to enhance coverage and explore unexplored paths. Santiago Balera and Junior (Balera and de Santiago Júnior, 2022) proposed a multi-perspective crawling approach using generative hyper-heuristic algorithms to capture diverse user behaviors and generate comprehensive test cases. Yousaf et al. (Yousaf et al., 2019) integrated model-driven techniques with crawling strategies to navigate interfaces and improve test case quality systematically. Jesus et al. (de Jesus et al., 2015) leveraged task-driven methods to generate navigation paths aligned with functional workflows. Nguyen et al. (Nguyen et al., 2022) applied Q-learning to explore dynamic state spaces, effectively detecting vulnerabilities like XSS. Pavanetto and Brambilla (Pavanetto and Brambilla, 2020) utilized deep learning models, combining RNNs and GANs to simulate realistic navigation paths, enhancing test coverage and interaction fidelity.
5.3.6. Template and Page Object-Based Testing
Robust and efficient test cases can be generated using templates or the Page Object Model. Yu et al. (Yu et al., 2015) used an incremental testing approach based on the Page Object Model that automatically generates page objects to represent Web pages. Neves et al. (de Almeida Neves et al., 2022) developed the Morpheus Web Testing tool, which extracts UI component information from JSF and Primefaces frameworks to generate system-level functional test cases, effectively covering complex user interaction scenarios. Leotta et al. (Leotta et al., 2015) utilized meta-heuristic techniques, including greedy and genetic algorithms, to automatically generate robust XPath locators, which reduces the maintenance overhead caused by DOM evolution.
5.3.7. Security Testing Generation Techniques
There are also techniques for generating test cases to detect security vulnerabilities in Web applications. Common vulnerabilities include for XSS and SQLi attacks. Akrout et al. (Akrout et al., 2014) proposed an automated black-box Web vulnerability identification and attack scenario generation method, enhancing vulnerability detection and accuracy. Awang et al. (Awang et al., 2019) used Cartesian product algorithms (Agesen, 1995) to generate SQLi test data, effectively detecting vulnerabilities. Wang et al. (Wang et al., 2016b) used UML statecharts to generate Web link security test scenarios, enhancing robustness in complex security environments. Nguyen et al. (Nguyen et al., 2022) used Q-learning algorithms (Jang et al., 2019), detecting security vulnerabilities, and demonstrating the potential for intelligent algorithms in Web security testing. Wang et al. (Wang et al., 2019) used evolutionary testing, combining client-side and server-side testing scenarios to coverage and efficiency, and demonstrating the potential for combinatorial optimization algorithms in Web security testing.
6. ANSWER TO RQ3: Test Case Execution
This section addresses RQ3, examining the execution of test cases within the context of WAT. The section explores the various methods, tools, platforms, environments, and key considerations of WAT testing: Each aspect is analyzed with a focus on its practical application, and its contribution to test-execution effectiveness.
6.1. Environments
The selection of testing environments is critical for WAT effectiveness, directly influencing the accuracy and scope of performance assessments. According to the literature, testing environments can be categorized into local (Wang et al., 2023), remote (Calzavara et al., 2021), and cloud-based (Wei et al., 2019).
Figure 5. Proportions of execution environments in surveyed papers.
Figure 5 shows that local environments were the most frequently used, accounting for 87.58%. The control and immediate feedback potential of local environments make them particularly suitable for iterative testing in early development (Everett and McLeod Jr, 2007). Remote environments (4.78%) support testing under various network conditions: This is essential for applications with geographically dispersed users. Cloud-based environments (7.64%) offer scalability and flexibility to meet complex testing demands and support continuous integration in large-scale applications.
6.1.1. Local Environments
The initial stages of development and testing often take place in local environments, which provide a controlled and isolated setting that facilitates rapid iteration and debugging. In these environments, because the web applications are deployed on the same machine that the tests are being executed on, the speed of iteration and bug fixing can be increased. Tools like Apache JMeter are increasingly adopted for testing web applications in local environments (Tiwari et al., 2023). JMeter excels in performance and functional testing and provides a reliable framework to validate isolated components during unit and integration testing, especially in controlled local settings.
A challenge for local environments, however, is that they cannot easily some simulate real-world scenarios like network latencies, distributed system behaviors, and large-scale performance (Prisadi et al., 2023). To address these limitations, tools like Docker have been widely used to create containerized environments that maintain consistency across different stages of development and production (Halin et al., 2019). Docker ensures that testing environments are uniform, thus reducing discrepancies among development, testing, and production setups (Marquardt and Buhl, 2021).
6.1.2. Remote Environments
Remote environments extend local capabilities by enabling the execution of test cases across multiple physical or virtual machines. Tools like Selenium Grid (Balera and de Santiago Júnior, 2022) play an important role in remote environments, enabling cross-browser and cross-platform testing (Debroy et al., 2018). Selenium Grid helps ensure that web applications function consistently across different environments and configurations (Jana and Oprea, 2019).
Despite their advantages, remote environments also face challenges, including network latency and the need to synchronize test execution across different systems (Zhou et al., 2014). Docker, in conjunction with Selenium Grid, helps alleviate these issues by ensuring that the test environments are consistent across remote nodes (Halin et al., 2019).
6.1.3. Cloud-Based Environments
Cloud-based environments offer the most flexible and scalable solution for large-scale WAT. Platforms such as Amazon Web Services (AWS), Google Cloud Platform (GCP), and Microsoft Azure offer dynamic resource allocation, allowing tests to be executed on-demand across multiple geographical locations (Ludinard et al., 2018; Su and Li, 2021). A cloud environment can simulate real-world scenarios (including high traffic and various network conditions), and can support global users (Arcuri et al., 2023). AWS Lambda, for example, is an important tool for cloud environments, providing a serverless architecture that allows testing to run without manual infrastructure management (Mukherjee et al., 2014). It can also dynamically expand resources based on test execution requirements, ensuring efficient test runs.
6.2. Execution Methods
Test-case execution has a critical role in the verification of web application functionality, security, and performance. Various methods can be used to execute test cases, each with certain advantages, depending on the nature and complexity of the application. This section provides a detailed exploration of the three primary approaches to test-case execution: manual, automated, and hybrid. Each approach is analyzed for its practical application, highlighting any contributions to the overall effectiveness of WAT test execution.
Figure 6. Proportions of execution methods in surveyed papers.
Figure 6 presents an overview of the s of different approaches to test-case execution in the surveyed literature. Automated execution dominates, appearing in 91.08% of the literature. This is as expected, given its efficiency and scalability for repetitive tasks and regression testing. Automated execution can support tasks requiring consistency across environments. Manual execution, although only in 1.91% of the surveyed papers, remains essential for tasks like exploratory testing and scenarios where human intuition is required. Hybrid execution, accounting for 7.01%, combines both approaches, optimizing test coverage and flexibility by automating routine tasks while leveraging manual testing for subjective analysis. This can be particularly valuable in large-scale applications.
6.2.1. Manual Execution
Manual execution involves human testers directly interacting with web applications to verify their behavior. This can be particularly effective for exploratory testing, as testers need flexibility to dynamically respond to application behavior and detect unexpected issues during testing. Exploratory testing can help to identify user-experience issues, or unexpected errors that are difficult to predict when designing automated scripts (Fan et al., 2023).
Manual testing is also often used to evaluate new features or complex user interfaces, which can require human judgment to assess usability, aesthetics, and overall user satisfaction (Leotta et al., 2020). However, there can be significant limitations to manual testing, including that it can be time-consuming, labor-intensive, and prone to human errors. This can be especially the case when dealing with large applications that require the execution of hundreds of test cases (Garg et al., 2015). Therefore, although manual testing remains essential for specific tasks, it may be most suitable for early development stages or areas that require subjective validation (Gupta and Bajpai, 2014).
6.2.2. Automated Execution
Automated execution is a more scalable and efficient approach, especially when repeating test cases in different environments or during regression testing. Typically, scripts and tools are used to execute test cases, capture results, and compare them with predefined expected outcomes (Leotta et al., 2020). This is often done without need for further human intervention. Selenium, for example, is widely used to automate functional testing across various browsers, ensuring consistent behavior and compatibility (Appelt et al., 2018).
Automated execution can significantly improve efficiency by reducing human involvement and allowing tests to run continuously in different environments. However, automated execution can require considerable upfront investment in scripting and tool configuration (Ammann and Offutt, 2017). Furthermore, it may not be suitable when subjective analysis is required, such as for visual-interface testing (Garg et al., 2015; Lounis et al., 2014).
6.2.3. Hybrid Execution
Hybrid execution combines both manual and automated techniques, drawing on the strengths of both approaches. In hybrid models, manual testing is typically used for exploratory tasks or validation of new features, while automated scripts handle repetitive tasks and regression testing  (Pelivani and Cico, 2021). This approach offers flexibility by allowing testers to manually address complex, subjective scenarios while taking advantage of automation for the repetitive, time-consuming processes (Guan et al., 2021; Garg et al., 2015).
It has been reported that combining UML activity diagrams with manual and automated test cases could improve overall test coverage and shorten execution time (Zhao and Gong, 2015). Similarly, a hybrid approach can be adopted in security testing, where potential vulnerabilities (such as SQLi points) can be detected automatically, and manual testers can then validate the findings (Singh et al., 2020). The combination of the thoroughness of manual testing and the efficiency of automation makes hybrid execution often the most practical solution for complex and large-scale web applications (Gupta and Bajpai, 2014).
7. ANSWER TO RQ4: Evaluations and Metrics
This section addresses RQ4 by exploring WAT metrics and evaluation methods. We first discuss the primary effectiveness metrics, such as test case quality, coverage, failure diagnosis, and vulnerability detection. This is followed by an analysis of efficiency metrics, such the time required for generation, execution, and error detection. Finally, empirical studies are reviewed to explain how these metrics have been applied to evaluate and improve WAT effectiveness and efficiency.
7.1. Test Effectiveness
Test effectiveness evaluates the ability of WAT methodologies to identify faults, vulnerabilities, and failures. This involves the use of specific metrics, including test-case quality, code coverage, failure-detection capability, and vulnerability-detection effectiveness. These metrics provide insights into the thoroughness and accuracy of WAT testing processes.
Figure 7. Proportions of test effectiveness metrics in surveyed papers.
Figure 7 presents an overview of the different test effectiveness metrics used in the surveyed literature. Failure-detection capability is the most frequently used metric, accounting for 30.22% of the papers. Code and coverage metrics (25.75%) measure the breadth of testing. Vulnerability detection comprises 20.90%, while test-case quality accounts for 16.42%. A small proportion of the literature, 6.72%, involved other metrics: For example, Kumar et al. (Saharan et al., 2018), and Tiwari et al. (Tiwari et al., 2023), for example, refer to performance-based metrics such as response time, error rate, and throughput.
7.1.1. Test-Case Quality
The quality of test cases directly affects the ability to detect defects in web applications. Dynamic testing approaches, such as Q-learning-based frameworks, have demonstrated substantial improvements in test-case adaptability. These methods optimize element interaction paths, increase branch and element coverage, and effectively identify critical defects in complex application environments (Fan et al., 2023). These tools provide a scalable solution for testing modern web applications by enhancing exploratory capabilities.
Prioritizing test cases based on risk and modification frequency has enhanced defect detection efficiency. These strategies uncover critical defects earlier by focusing on high-risk and frequently modified components, which streamlines testing processes and improves efficiency and effectiveness (Ali-Shahid and Sulaiman, 2015).
Furthermore, fast evaluation techniques that leverage clustering-based sampling can improve test-case efficiency in large-scale applications. These methods group similar application paths and ensure representative coverage, which significantly reduces redundant testing efforts while maintaining reliability (Zhang et al., 2015b).
7.1.2. Code Coverage Metrics
Code coverage metrics — including line, branch, and function coverage — can help show how thoroughly an application’s codebase is exercised during testing (Nguyen et al., 2019; Wang et al., 2016a). Coverage analysis tools like Selenium allow testers to capture interactions across multiple layers of the web application, from the front-end UI to the back-end logic (Paul and Tommy, 2018).
However, high coverage alone does not guarantee comprehensive fault detection. Research has shown that combining coverage metrics with fault injection methods enhances the ability to uncover defects (Chaleshtari et al., 2023). Injecting controlled faults into the system, and observing whether or not the test cases can detect them, provides a more rigorous assessment of a test suite’s fault-detection capabilities (Gupta and Bajpai, 2014). Mutation testing complements code coverage metrics by introducing controlled changes (i.e., “mutants”) to evaluate the fault-detection capabilities of test cases (Habibi and Mirian-Hosseinabadi, 2015; Leotta et al., 2024). While coverage metrics measure the breadth of testing, mutation testing ensures that covered paths are rigorously tested, enhancing the reliability of the testing process.
7.1.3. Failure-Detection Capability
Failure-detection capability refers to the accurate identification and localization of faults during testing. Techniques such as dynamic trace-based analysis compare the execution traces of different versions of the application to identify discrepancies (Touseef et al., 2019). This approach has proven effective in failure detection, particularly when combined with automated tools for analyzing execution logs and detecting potential issues, such as security vulnerabilities, in large-scale applications. This can help testers to pinpoint the specific lines of code or configuration changes responsible for faults, significantly reducing the time required for debugging and resolution (Walsh et al., 2015; Halin et al., 2019).
Automated diagnostic tools further improve this process by analyzing execution logs and error traces to identify the exact points of failure (Sunman et al., 2022). These tools are particularly useful for large-scale web applications, where manual fault-localization could be time-consuming and error-prone (Fan et al., 2023).
7.1.4. Vulnerability Detection
Vulnerability detection includes the identification of potential security issues such as SQLi and XSS that could be exploited by attackers (Awang et al., 2019; Mokbal et al., 2019). Automated vulnerability detection tools simulate real-world attack scenarios, testing the robustness of the application’s security defenses (Nagpure and Kurkure, 2017).
Recent advances in machine learning have enhanced vulnerability-detection capabilities. These systems analyze traffic and behavioral patterns to identify anomalies that may indicate vulnerabilities (Jana and Oprea, 2019). As machine learning models learn from new data, they continuously improve their ability to detect both known and unknown threats, providing a dynamic defense against emerging security risks (Calzavara et al., 2020).
7.2. Test Efficiency
Test efficiency is a core WAT metric, directly affecting the cost effectiveness of the testing process. Efficiency metrics assess the resources required for comprehensive test coverage and fast fault detection. Some of the main metrics include the time taken for test case generation (Gupta and Bajpai, 2014) and execution (Pelivani and Cico, 2021), and error detection (Fan et al., 2023).
Figure 8. Proportions of test efficiency metrics in survey papers.
Figure 8 presents an overview of the different test efficiency metrics used in the surveyed literature. Test execution time and test case generation time are the most frequently measured metrics, each accounting for 30.83% of the literature. Error detection time is slightly more widely used, appearing in 32.33% of the surveyed papers. A smaller proportion, 6.02%, includes other metrics: For instance, Sung et al. (Sung et al., 2016) used static analysis methods to reduce redundant tests, thereby improving testing speed and efficiency.
7.2.1. Test Case Generation Time
The test-case generation time measures how efficiently test cases are created and is a major part of the overall testing efficiency. Several techniques have been developed to streamline this process. The Page Object Pattern (POP) method has also been able to reduce the test-case generation time as the test suite grows due to the reuse of defined objects. This approach improves efficiency as the same components are used across multiple test cases (Leotta et al., 2020). Studies involving WebQT, which employs machine learning algorithms, have shown that test-case generation can be significantly accelerated through optimized exploration and test-sequence generation (Gupta and Bajpai, 2014). Tools like CRaTCP can automatically generate and execute all possible test case combinations to further reduce the manual efforts (Nagowah and Kora-Ramiah, 2017).
7.2.2. Test-Execution Time
The test-execution time refers to the time taken to execute the generated test cases. Automated testing frameworks like Selenium and Katalon Studio can dramatically reduce execution time (compared to manual testing) (Pelivani and Cico, 2021). Research has confirmed their effectiveness in optimizing test execution times, particularly in large-scale web applications (Pelivani and Cico, 2021). Genetic algorithms, for example, have been used to reduce execution time by optimizing test paths, leading to earlier detection of faults (Stallenberg and Panichella, 2019). Concurrent execution techniques, where multiple test cases are run in parallel, can also further reduce execution times (Garg et al., 2015). Prioritization techniques that execute high-risk test cases first can lead to more efficient use of resources by focusing on the most critical components.
7.2.3. Error-Detection Time
The error-detection time reflects how quickly errors or vulnerabilities are identified during testing. Hybrid methods that integrate static and dynamic analysis improve detection efficiency by focusing on critical code paths and minimizing redundant analysis (Zhao and Gong, 2015). Techniques leveraging anomaly detection through multi-model approaches accelerate fault detection by identifying deviations in real-time and ensuring low false positive rates (Zhang et al., 2017). By minimizing manual intervention, automated test suites enable real-time fault identification and significantly accelerate the detection process (Paul and Tommy, 2018). Moreover, combining diverse static analysis tools has significantly improved detection speed and accuracy, particularly in scenarios with limited resources or time constraints (Nunes et al., 2019).
7.3. Empirical Studies
This section examines the use of empirical studies in WAT, focusing on effectiveness and efficiency metrics. It explains how these studies can enhance testing performance, optimize testing processes, and offer insights for further improvements in testing tools.
7.3.1. Application of Effectiveness and Efficiency Metrics in WAT
Coverage has been widely used to measure testing effectiveness in empirical WAT studies (Garousi et al., 2021; Mattiello and Endo, 2022). MBT can generate high-quality test cases, with good code coverage and fault-detection capabilities (Nabuco and Paiva, 2014; Nguyen et al., 2019). Compared to manual testing, MBT provides a more systematic approach to cover application requirements, especially for complex systems, and achieves better testing efficiency (Wang et al., 2019). MBT is particularly effective at handling the complex interactions between multiple modules or interfaces within large-scale web applications. Empirical studies have shown that MBT reduces test redundancy and optimizes test case generation (Hanna and Munro, 2018).
Automated tools like Model-Based Testing Leveraged for Web Tests (MoLeWe) (Mattiello and Endo, 2022) have demonstrated excellent testing efficiency (Li and Zeng, 2014): MoLeWe can generate high-quality test cases, increasing test coverage while reducing execution time (Garousi et al., 2021). It can significantly reduce manual testing time, especially in applications with frequent version updates. MoLeWe has also demonstrated flexibility in resource allocation, optimizing system resource utilization in multi-threaded testing.
Fault Detection Density (FDD) and Fault Detection Effectiveness (FDE) are core effectiveness metrics used to evaluate the fault-detection capabilities of testing tools (Habibi and Mirian-Hosseinabadi, 2015). The MUTANDIS mutation-testing tool can generate non-equivalent mutants, significantly improving the fault-detection rate in JavaScript applications (Mirshokraie et al., 2014). Mutation testing can covers a wide range of fault scenarios using diverse mutants: This can improve the comprehensiveness of the test cases. MUTANDIS has also been shown to enhance fault-detection efficiency while reducing resource waste caused by redundant testing paths.
Efficiency metrics are not only used for coverage and fault-detection capabilities, but also in execution time and resource utilization (Putri et al., 2017; Ali and Badr, 2015). Empirical studies have shown that tools like JMeter can optimize performance testing through distributed architectures (Saharan et al., 2018; Kiran et al., 2015). By parallelizing the execution of test cases, JMeter can significantly reduce execution time while ensuring system stability under high-load conditions. This architecture has been shown to perform well in real-world scenarios, especially in large-scale multi-user concurrent environments (Guarnieri et al., 2017).
7.3.2. Improvements in Testing Performance through Empirical Studies
Empirical studies have confirmed the effectiveness of reinforcement learning techniques in the prioritization of test cases (Chang et al., 2023): High-risk areas can be prioritized, thereby reducing overall testing time and improving fault-detection efficiency.
Automated tools can also improve testing efficiency. MoLeWe can automatically generate test cases and scripts, minimizing manual intervention and significantly enhancing automation (Mattiello and Endo, 2022): This not only reduces the error rate from manual operations but also enables the rapid adaptation of test suites to accommodate new features or updates.
8. ANSWER TO RQ5: Available Tools
This section addresses RQ5, identifying and cataloging the WAT tools that are currently available. This should provide an overview of the technological resources available to support this field. We first list the tools, and then classify them based on their types and features. The section concludes with some details on tool accessibility, including links for further exploration.
8.1. Testing Tools
Various WAT tools are available to streamline and automate the testing process, supporting efficiency, scalability, and consistency across diverse environments. These tools can support the automation of various different types of testing (including functional, regression, security, and performance testing), helping to examine the reliability and robustness of web applications.
Figure 9 presents an overview of the proportions of testing tools found in the surveyed literature. Security testing tools account for 41.57%, highlighting the increasing focus on mitigating cybersecurity risks. Functional and regression testing tools represent 38.95%, emphasizing their critical role in validating application functionality and maintaining system stability following changes. Performance testing tools comprise 11.34%, and cross-platform/distributed environment tools make up 8.14%, reflecting their specialized contributions to optimizing system performance and ensuring consistency across platforms. This distribution reveals varying priorities in testing, with security and functional testing receiving the most attention, and performance and cross-platform testing appearing to be applied only as required.
Figure 9. Proportions of testing tools in surveyed papers.
8.1.1. Functional and Regression Testing Tools
Functional and regression testing tools, like Selenium (Pelivani and Cico, 2021) and Katalon Studio (Pelivani and Cico, 2021), have been widely used for cross-browser and cross-platform testing: By automating test execution across various browsers and operating systems, they help to verify that web applications function correctly after updates and code changes
8.1.2. Security Testing Tools
Tools like Burp Suite (Nagpure and Kurkure, 2017) and OWASP ZAP (Amankwah et al., 2020) can help to identify security vulnerabilities, including SQLi, XSS, and other common web application threats. These tools are essential for protecting web applications from attacks, and ensuring compliance with security standards.
8.1.3. Performance Testing Tools
JMeter (Fan et al., 2023) and LoadRunner (Khan and Amjad, 2016) have been used to simulate high-traffic loads, helping to evaluate how web applications perform under stress. They assess critical aspects of performance, including response times and throughput, evaluating whether or not the application remains stable under heavy load conditions.
8.1.4. Cross-platform and Distributed Environment Tools
Tools such as Docker (Marquardt and Buhl, 2021) and Selenium Grid (Kiranagi and Shyam, 2017) enable test cases to be executed in distributed environments and containerized platforms. They help to ensure consistency in execution and provide scalability for larger systems, supporting the efficient management of complex testing scenarios.
8.2. Tool Accessibility
The ease of access to WAT tools is a crucial factor in their adoption, and eventual effectiveness in the software development and testing community. Table LABEL:TAB:wat_tools lists some popular WAT tools [^1] To measure the popularity of each WAT tool, we collected the number of times to be mentioned in the selected 314 primary studies., categorized by type, core functionalities, and access links: This structured presentation facilitates the selection of tools based on specific testing requirements, ranging from performance evaluation and security vulnerability detection to automated UI testing and code-quality analysis. Due to page limit constraints, we provide a comprehensive summary of nearly all relevant tools, available in the online supplementary material (Li et al., 2024b).
Table 4. A comprehensive overview of tools.
Tool Name Tool Type Overview Resource Link Reference
Selenium Open-source It is open source, cross-platform, and supports multiple languages and browsers. https://www.selenium.dev/
https://github.com/SeleniumHQ
 (Aditya et al., 2023)
ZAP Open-source It is an open-source web application scanner that excels at finding common web application vulnerabilities. https://www.zaproxy.org/
https://github.com/zaproxy
 (Amankwah et al., 2020)
JMeter Open-source It is an open-source tool for load testing and performance measurement, especially for web applications. It allows the simulation of a large number of users and provides detailed reports. https://jmeter.apache.org/
https://github.com/apache/jmeter.
 (Cai and Hu, 2014)
Arachni Open-source It is a web vulnerability scanner designed to automatically detect and assess security vulnerabilities in web applications. https://github.com/Arachni/arachni  (Román Muñoz et al., 2018)
Crawljax Open-source It is a tool for automatically crawling and testing web applications, which automatically explores the GUI of a web application, creates state flow diagrams, and generates test cases based on interactions. https://github.com/crawljax/crawljax  (Sunman et al., 2022)
Cucumber Open-source It is a tool that automatically runs business case-driven tests, ensuring that all scenarios are tested as per the acceptance criteria. https://github.com/cucumber/cucumber  (Rathod et al., 2015)
Nikto Open-source It is an open-source command line tool that focuses on quickly scanning web server and website configuration errors, known security vulnerabilities, and missing XSS protection headers. https://cirt.net/Nikto2  (Devi and Kumar, 2020)
Puppeteer Open-source It automates web interactions to perform testing. https://github.com/puppeteer/puppeteer  (Chen et al., 2021)
Cypress Open-source It is a JavaScript-based tool for automated end-to-end testing of modern web applications. https://github.com/cypress-io/cypress  (Mollah and van den Bos, 2023)
Axiom Open-source It is a dynamic infrastructure framework for distributed computing and testing that automates distributed penetration testing across cloud environments. https://github.com/pry0cc/axiom  (Prisadi et al., 2023)
Beautiful Soup Open-source It can be used to extract data from html and xml files. https://www.crummy.com/software/BeautifulSoup
https://git.launchpad.net/beautifulsoup
 (Paul and Tommy, 2018)
Commix Open-source It is a tool that automatically detects and exploits command injection vulnerabilities in web applications, including classic command injection, blind command injection, and other types of tools. https://github.com/commixproject/commix  (Stasinopoulos et al., 2019)
Katalon Studio Commercial (Free) It is easy to install and use, has powerful built-in reporting capabilities, and supports record and playback. https://katalon.com/  (Nguyen et al., 2021)
Cytestion Open-source It is a web-based tool for automated exploration and testing of GUIs. https://gitlab.com/lsi-ufcg/cytestion/cytestion  (Moura et al., 2023)
Lazyrecon Open-source It is an automated vulnerability scanning and data collection framework. https://github.com/nahamsec/lazyrecon/blob/master/lazyrecon.sh  (Shaji and Subramanian, 2021)
Jena API Open-source It is an open-source Java framework for building semantic web and linked data applications, supporting SPARQL queries. https://jena.apache.org/  (Durai et al., 2021)
Binwalk Open-source It is a firmware analysis tool suitable for most firmware unpacking. https://github.com/ReFirmLabs/binwalk  (Li et al., 2023)
jÄk Open-source It is a web application crawler and scanner tool that uses dynamic Java Script code analysis. https://github.com/ConstantinT/jAEk  (Pellegrino et al., 2015)
Skipfish Open-source It is an efficient web application security vulnerability scanning tool that can quickly detect common web application security issues. https://code.google.com/archive/p/skipfish/  (Touseef et al., 2019)
Sqlmap Open-source It is a tool for detecting and exploiting SQL injection vulnerabilities. https://sqlmap.org/  (Prisadi et al., 2023)
Scrapy Open-source It is a tool for quickly crawling HTML-based static web page content. https://github.com/scrapy/scrapy  (Rennhard et al., 2022)
Metasploit Open-source It is an open-source framework for penetration testing and vulnerability exploitation. https://www.metasploit.com/
https://github.com/rapid7/metasploit-framework
 (Jana and Oprea, 2019)
OWASP AppSensor Open-source It is a tool that detects suspicious application layer behavior and triggers events based on predefined sensors. https://owasp.org/www-project-appsensor/  (Shahrivar et al., 2023)
HTMLCS Open-source It is a tool for checking whether HTML code complies with WCAG standards. http://squizlabs.github.io/HTML_CodeSniffer/  (Campoverde-Molina et al., 2021)
GPTFuzzer Open-source It is a tool for generating web application firewall (WAF) bypass payloads, testing and discovering common web attacks such as SQL injection, cross-site scripting (XSS), and remote command execution (RCE) vulnerabilities. https://github.com/hongliangliang/gptfuzzer  (Liang et al., 2023)
SideeX Open-source It is a tool that provides cross-browser Web UI automated testing, supports recording and playback functions, and optimizes the test process through an automatic waiting mechanism. http://sideex.org  (Lee et al., 2018)
CICFlowMeter Open-source It is a tool that provides network flow signatures for intrusion detection and malicious behavior classification. https://github.com/ahlashkari/CICFlowMeter  (Rajić et al., 2023)
OWASP AppSensor Open-source It is a tool that detects suspicious application layer behavior and triggers events based on predefined sensors. https://owasp.org/www-project-appsensor/  (Shahrivar et al., 2023)
WebPageTest Commercial (Free) It is used to analyze web page performance, providing details such as page load time, resource load time, etc. https://www.webpagetest.org/  (Ali-Shahid and Sulaiman, 2015)
WebTest Open-source It is a tool that supports visualization of output coverage, helps testers understand covered and uncovered output parts, and optimizes test cases. https://github.com/git1997/VarAnalysis  (Nguyen et al., 2019)
GTmetrix Commercial (Free) It is used to measure web page loading speed and performance and provide detailed optimization suggestions. https://gtmetrix.com/  (Ali-Shahid and Sulaiman, 2015)
SikuliX Open-source It is an image-based testing tool that can control user input methods to interact with GUI elements, suitable for both web and non-web applications. http://sikulix.com  (Macchi et al., 2021)
Hakrawler Open-source It is a web crawler tool that collects web links and identifies injection points. https://github.com/hakluke/hakrawler  (Prisadi et al., 2023)
Axe Studio Commercial (Free) It is a tool that focuses on automated detection of web accessibility issues. https://www.deque.com/axe/  (Campoverde-Molina et al., 2021)
Cornipickle Open-source It is a declarative testing tool that allows developers to define layout errors using logical expressions and compare states of different snapshots. https://github.com/liflab/cornipickle  (Beroual et al., 2020)
Open-source tools like Selenium (Dawei et al., 2016), JMeter (Cai and Hu, 2014), and ZAP (Touseef et al., 2019) provide cross-platform compatibility, support multiple programming languages, and have received strong community support. They benefit from scalability, flexibility, and collaborative development through platforms like GitHub, where contributors can provide updates and improvements.
In contrast, commercial tools such as Katalon Studio (Nguyen et al., 2021) and GTmetrix (Ali-Shahid and Sulaiman, 2015) often prioritize enhanced usability and integrated reporting features, making them more suitable for testers and organizations that value refined interfaces and out-of-the-box functionality. However, limitations in the free versions may make them less suitable for certain users.
Nearly all of the reviewed tools are easily accessible through official websites or open-source platforms. This enables a smooth integration into a wide range of testing environments. Comprehensive documentation and active community support further simplify adoption and customization. The tools offer a range of diverse capabilities: From identifying security vulnerabilities with tools like Wapiti (Anagandula and Zavarsky, 2020) and Skipfish (Touseef et al., 2019), to enhancing page performance with GTmetrix (Ali-Shahid and Sulaiman, 2015) and WebPageTest (Ali-Shahid and Sulaiman, 2015). Their wide availability and applicability underscore their role in advancing WAT practices.
9. ANSWER TO RQ6: Challenges and Future Work
This section addresses RQ6, identifying unresolved challenges in WAT, and suggesting directions for future research. We start by outlining the key challenges still affecting the field. We then discuss potential areas for future work that could help address these issues and improve WAT practices.
9.1. Challenges
Despite notable advances over recent years, a number of significant challenges to WAT remain that prevent the full realization of its potential. Here we discuss several key challenges:
  • As web applications increase in complexity, test automation faces significant challenges, particularly when scaling up to handle large test suites efficiently.
    The recent use of machine learning and containerization to manage these large-scale testing processes has helped. However, limitations persist for dynamic content and asynchronous operations (Pellegrino et al., 2015). Additionally, tool fragmentation across platforms continues to affect scalability (Pelivani and Cico, 2021).
  • Balancing test coverage and execution efficiency remain challenges in web testing. Virtual DOM coverage techniques have demonstrated the potential to enhance testing efficiency for dynamic web applications by optimizing coverage metrics. However, achieving a balance between efficiency and comprehensive coverage remains an open challenge (Zou et al., 2014). Hyper-heuristic algorithms can reduce test-case redundancy, but more research into their application is required (Balera and de Santiago Júnior, 2022).
  • Failure diagnosis can be a bottleneck for WAT. Model-based approaches can help, but false positives and false negatives remain significant issues (Alkofahi et al., 2022). Furthermore, automated repair strategies may have difficulties with the dynamic content in complex web applications, indicating a need for more robust tools.
  • Maintaining test suites for dynamic web applications can also present significant challenges: Frequent updates to UI and DOM structures often break test cases (Ryou and Ryu, 2018). Although recent work has used machine learning to adapt test scripts to changes, this remains limited when dealing with significant UI overhauls.
  • Web testing tools remain fragmented, leading to inefficiencies in end-to-end testing processes. Tools like Selenium and Katalon may excel in specific areas, but they also struggle to provide cohesive solutions for more comprehensive testing needs. Recent studies have emphasized the importance of integrating these tools for more seamless cross-platform testing (Marquardt and Buhl, 2021). However, ensuring that these tools can work together across different testing environments remains a challenge (Pelivani and Cico, 2021).
  • The lack of standardized metrics for evaluating testing effectiveness remains a significant challenge. Although various tools provide performance and security-testing capabilities, they often lack consistent benchmarks, making comparisons difficult. There is a need for unified evaluation frameworks that provide reliable metrics for test coverage and efficiency (Sönmez and Kiliç, 2021). Initiatives like WebEV are helping to standardize some aspects of test-behavior analysis, but further research is necessary  (Fuad and Sakib, 2023).
9.2. Future Work
This section outlines some potential future WAT research directions, including enhancing scalability, the integration of Large Language Models (LLMs), enhanced failure diagnosis, and the development of standardized evaluation metrics.
  • Future research should enhance the scalability of test-automation frameworks to accommodate the growing complexity of modern web applications. Machine learning models should be further optimized, reducing redundancy while maintaining coverage. The use of containerization technologies, like Docker and Kubernetes, should be expanded to better enable parallel execution and to improve resource efficiency. Integrating tools into unified frameworks will also enhance automation processes and reduce fragmentation (Pelivani and Cico, 2021).
  • LLMs represent a promising approach for automating test-case generation, particularly for web forms where contextual information is key. Initial studies (Yao et al., 2024) have shown the potential of models like GPT-4 (Sanderson, 2023) and Baichuan2 (Xiao et al., 2024). However, further research is required to address security concerns and refine the models for real-world use, including for dynamic content (Li et al., 2024a).
  • Failure-diagnosis tools are critical to the precision and reliability of WAT. Many current systems have difficulties with false positives and false negatives: Future research should therefore explore adaptive systems that can better interpret the complex interactions in web applications. Combining functional and cross-browser testing tools will also enhance overall efficiency (Alkofahi et al., 2022; Pelivani and Cico, 2021).
  • Standardized metrics for WAT-tool effectiveness are also needed. The absence of universal evaluation frameworks for coverage, efficiency, and performance limits meaningful comparisons: Consistent metrics will enable more accurate assessments and foster innovation in WAT methodologies (Sönmez and Kiliç, 2021; Fuad and Sakib, 2023).
  • Multi-agent reinforcement learning (MARL) has also shown potential for enhancing test coverage and efficiency in complex web applications  (Fan et al., 2024). Initial studies have demonstrated MARL’s effectiveness for single-agent approaches. However, further research is needed to refine inter-agent communication and data-sharing protocols, reducing redundancy and optimizing exploration across web states.
  • Large vision-language models (LVLMs), such as VETL (Wang et al., 2024), can support context-aware GUI testing by generating relevant inputs and selecting UI elements based on visual context. Although LVLMs have shown potential in managing dynamic, visually complex interfaces, future work should explore hybrid approaches that combine LVLMs with heuristic methods to improve cost-efficiency and adaptability, especially in applications with frequent content updates.
10. Conclusions
This review paper has offered a detailed analysis of advances in WAT over the past decade. The review has examined a collection of computer science papers published between January 1, 2014, and December 31, 2023. A rigorous review-research methodology was employed: Five major academic databases were systematically searched, using targeted keywords, with each retrieved paper carefully examined to ensure its relevance to the study.
Our review was guided by eight research questions, which helped organize and synthesize the data. These questions addressed essential aspects of WAT, including test-case generation and execution, evaluation metrics, and available tools. The paper provides a clear understanding of the key trends and challenges in WAT over the past decade.
Various test-case generation methods were reviewed, including those designed to handle the dynamic and asynchronous nature of modern web applications. Regarding test execution, the importance of testing across different environments to reflect real-world conditions was emphasized. Besides, metrics used to evaluate WAT effectiveness, such as test coverage and error-detection rates, were also explored, highlighting the testing process’s efficiency and quality. Additionally, available WAT tools were categorized based on their functions and accessibility, providing a practical overview for researchers and practitioners.
This paper represents a useful resource for researchers and practitioners aiming to enhance the quality, reliability, and security of web applications.