### v2

# A Survey on Web Application Testing: A Decade of Evolution

Abstract.
As one of the most popular software applications, a web application is a program, accessible through the web, to dynamically generate content based on user interactions or contextual data, for example, online shopping platforms, social networking sites, and financial services. Web applications operate in diverse environments and leverage web technologies such as HTML, CSS, JavaScript, and Ajax, often incorporating features like asynchronous operations to enhance user experience. Due to the increasing user and popularity of web applications, approaches to their quality have become increasingly important. Web Application Testing (WAT) plays a vital role in ensuring web applications’ functionality, security, and reliability. Given the speed with which web technologies are evolving, WAT is especially important. Over the last decade, various WAT approaches have been developed. The diversity of approaches reflects the many aspects of web applications, such as dynamic content, asynchronous operations, and diverse user environments. This paper provides a comprehensive overview of the main achievements during the past decade: It examines the main steps involved in WAT, including test-case generation and execution, and evaluation and assessment. The currently available tools for WAT are also examined. The paper also discusses some open research challenges and potential future WAT work.
Software testing, web application testing, survey
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software verification and validationccs: Information systems Web applicationsccs: General and reference Surveys and overviews
1. Introduction
The Software Development Life Cycle (SDLC) (Shafiq et al., 2021) is composed of several essential phases, including requirements specification, architecture design, coding, testing, and maintenance (Hanna and Ahmad, 2022). Each phase involves specific processes and activities that contribute to the overall development of the software. Among these phases, testing is particularly critical, often accounting for a significant proportion of time and cost for developing a software application. Recent studies (López-Martín, 2022; Bindu Bhargavi and Suma, 2022) indicate that the testing process can consume a range of 10% to 80% of total development time, depending on the complexity of the project. The primary objective of testing is to identify errors that manifest during program execution. An effective testing methodology improves the probability of detecting such errors, for example, by systematically running and examining a wide range of execution paths, including edge cases and fault-prone areas (Li et al., 2014b; Mahdieh et al., 2022).
Since the adoption of the World Wide Web in the early 1990s (Berners-Lee, 1992), web applications have become integral to various sectors, including finance, healthcare, education, and commerce (Doğan et al., 2014). These applications have introduced innovative features and functionalities that facilitate consistent and efficient access to information across diverse sources, significantly transforming how society interacts with digital content (Jazayeri, 2007). Meanwhile, web applications rely on a set of core technologies, including HTML to define the structure of web pages, CSS to style, JavaScript to enable interactivity, and Ajax for asynchronous data exchanges between the client and the server. These technologies are the foundation of modern web applications and are critical in supporting their dynamic and distributed nature (Garrett et al., 2005; Arora and Sinha, 2013; Sampath and Sprenkle, 2016).
As web applications grow increasingly complex and technologies continue to evolve, the requirement for robust and effective testing methodologies has become more crucial atkinson2002towards: It is necessary to continuously evolve and advance testing strategies to ensure the reliability, security, and overall quality of web applications (Li et al., 2014b). Web Application Testing (WAT) (Doğan et al., 2014) has become an essential process in the SDLC of web applications, which is used to evaluate and ensure the quality of Applications Under Test (AUTs) (Karimoddini et al., 2022), making it a cornerstone of modern software quality assurance. WAT aims to ensure that web applications function as intended, meet specified requirements, and are free from vulnerabilities. The scope of WAT goes beyond bug detection and fixing, and includes enhancing user experience, ensuring data security, and maintaining the overall reliability of the web systems. As web technologies evolve, testing methodologies have advanced to address web applications’ increasing complexities and demands.
Web applications have had an important impact on society (Garousi et al., 2013), reshaping various aspects of daily life and industry (Onukrane et al., 2023). Despite their significance, however, WAT reviews have tended to be narrowly focused (Prazina et al., 2023), addressing specific aspects without providing a comprehensive overview (Kousar et al., 2023). One WAT survey paper, by Prazina et al. (Prazina et al., 2023), only concentrates on the testing and analysis methods for automated web layouts. In contrast, significant attention has been devoted to the security aspects of web applications (Rahman and Izurieta, 2022; Aydos et al., 2022; Zhang et al., 2021; Altayaran and Elmedany, 2021; Cui et al., 2020; Wang et al., 2020; Muslihi and Alghazzawi, 2020; Marashdih et al., 2019; Kumar and Sheth, 2016; Rafique et al., 2015; Gupta et al., 2014). Additionally, Gupta et al. (Gupta et al., 2018) focused on the automated generation of regression test cases for web applications.
Early reviews of WAT, such as those by Li et al. (Li et al., 2014b) and Dougan et al. (Doğan et al., 2014), established foundational insights into testing dynamic web environments, focusing on early challenges and techniques. More recently, Balsam and Mishra (Balsam and Mishra, 2024) have provided a review to discuss some challenges and opportunities in WAT. However, this paper focused on only 72 primary studies, which means that many significant contributions and broader advancements, over the past decade, have been overlooked. To address this gap, we aim to provide a comprehensive examination of the latest progress in WAT, detailing core processes, evaluating current tools, and exploring emerging research directions to offer a thorough overview of the field’s evolution over the past decade.
This paper provides a comprehensive examination and discussion of WAT across its entire life cycle. We explore the research status and WAT advances from the perspectives of test case generation and execution, failure diagnosis, evaluation and assessment, regression testing, and available tools. We also examine some key open research challenges and discuss potential directions for future WAT work.
The rest of this paper is organized as follows: Section 2 provides background information on the formal definition of WAT and the fundamental testing techniques employed in the field. Section 3 outlines the survey methodology, including presenting the six research questions, the literature retrieval process, and the statistical analysis of the retrieved studies. Sections 4 to LABEL:SEC:rq8 address each of the research questions (RQ1 to RQ6), respectively. Finally, Section 10 concludes the paper by summarizing key findings and proposing directions for future research.
2. Background
This section presents some preliminary WAT concepts and provides an overview of web applications and WAT, covering some core definitions and relevant testing techniques.
2.1. Web Applications
A web application (Jazayeri, 2007) is a software system that facilitates user interaction through a combination of front-end and back-end components, over a network, typically accessed through a web browser. The front-end, or client-side, renders the user interface and handles user inputs. It consists of web pages constructed using technologies such as HTML, CSS, and JavaScript. This layer is designed to provide a responsive and interactive user experience, enabling seamless communication between the user and the application. The back-end, or server-side, encompasses the core functionalities of the application: These include data processing, business logic execution, and database management. The back-end is the key to the application: It processes requests from the front-end and ensures that the correct data and functionalities are provided to the user. Communication between the front-end and back-end typically relies on protocols such as HTTP and HTTPS, which support smooth data exchange and maintain the application’s state (Kumar et al., 2017; Neumann et al., 2018). Together, these components provide an integrated system that supports a wide range of user interactions, from simple content browsing to complex transaction processing.
Web applications can be either static or dynamic. Static web applications have fixed content that remains unchanged regardless of user interactions, making them suitable for delivering straightforward information, such as in digital brochures or basic company websites (Vaidyanathan and Mautone, 2009). Dynamic web applications, in contrast, provide a more interactive experience, with content that adapts to user inputs, interactions, or real-time data (Sampath and Sprenkle, 2016). The dynamic nature of these applications allows for greater flexibility and user engagement, making them well-suited for complex platforms like social media networks, e-commerce sites, and online service portals that require continuous interaction and real-time responsiveness (Fulcini and Ardito, 2022).
2.2. Web Application Testing
Web applications have become a fundamental part of our daily lives, making thorough testing a vital process to ensure their functionality and dependability. Effective testing ensures that these applications not only meet the required functionality but also have sufficient security, and good performance and reliability. Thorough testing of web applications can help to identify and address potential vulnerabilities, performance bottlenecks, and functional issues: This enhances the user experience and helps to safeguard sensitive data, which is vital for maintaining trust in web-based services. Comprehensive testing ultimately contributes to the development of robust, dependable web applications that can adapt to the dynamic demands of users and the evolving technological landscape.
WAT involves several key steps:
  • Analysis of the functional and non-functional requirements of the web application to guide the development of comprehensive test cases.
  • Generation of test data according to the testing methodology (such as model-based or security testing), and appropriate design of test cases, to ensure a broad coverage of potential scenarios.
  • Execution of the test cases by sending HTTP requests to the web application, and comparison of the responses with the expected outcomes to identify inconsistencies.
  • Documentation of the test results, analysis of any issues, and application of regression testing to ensure that recent changes have not introduced new defects.
  • Implementation of continuous testing and monitoring to maintain the application’s stability and security as it evolves.
3. Methodology
In this section, we present the six research questions (RQs) related to WAT that structured our study, and this paper. Our WAT survey involved a systematic and structured methodology, guided by the published frameworks of Huang et al. (Huang et al., 2021) and Webster et al. (Webster and Watson, 2002). Our findings’ presentation draws from recent surveys on related topics, including the testing of RESTful APIs (Golmohammadi et al., 2023), mutation analysis (Jia and Harman, 2010), and metamorphic testing (Segura et al., 2016; Chen et al., 2018). The following sections detail key aspects of the review process and results.
3.1. Research Questions
The goal of this survey paper was to identify and categorize the available WAT information. To achieve this, the following six RQs guided the study:
  • RQ1: How has WAT evolved, and what WAT topics have been examined in published studies?
  • RQ2: What methods and approaches have been used for WAT test case generation?
  • RQ3: How are test cases executed in WAT?
  • RQ4: What metrics are used to evaluate WAT? How are they used?
  • RQ5: What tools are available to support WAT?
  • RQ6: What are the challenges and future work related to WAT?
An answer to RQ1 should provide an understanding of the WAT literature and the distribution of its key topics. RQ2 prompted an examination of the various methods and approaches used for WAT test case generation, revealing the application of a diversity of techniques. RQ3 focuses on the execution of WAT test cases, aiming to clarify the processes and practices involved. RQ4 should reveal the metrics and evaluation frameworks used in WAT, enabling an understanding of how the WAT effectiveness and efficiency are assessed. RQ5 should reveal the currently available WAT tools, offering a comprehensive overview of the technological resources for WAT. Finally, based on the insights gathered from the other RQs, RQ6 shall identify unresolved challenges and propose potential directions for future WAT research. Figure 1 presents an overview of the structure of this paper.
Figure 1. Structure of this survey paper.
3.2. Literature Search and Selection
Similar to previous survey studies (Huang et al., 2021; Webster and Watson, 2002; Golmohammadi et al., 2023; Harman et al., 2015; Anand et al., 2013), we selected the following online repositories of technical research literature:
  • ACM Digital Library (ACM)
  • Elsevier Science Direct (Elsevier)
  • IEEE Xplore Digital Library (IEEE)
  • Springer Online Library (Springer)
  • Wiley Online Library (Wiley)
These repositories were chosen for their extensive collections of conference, symposium, and workshop papers, as well as for their access to multiple journals that are considered highly relevant to WAT (Huang et al., 2021). Our survey examined computer science papers published between January 1, 2014, and December 31, 2023.
Once the literature repositories were identified, search strategies were developed and tailored for each of them, using WAT-specific terminology and formatting. Given the unique characteristics and limitations of the advanced search functions in each repository (such as Elsevier’s restriction on using a maximum of eight Boolean connectors per field), the search methods were designed to strictly adhere to the repository rules and constraints. To ensure comprehensive coverage of all relevant literature, multi-dimensional and multi-layered filtering keywords were used to maximize the retrieval of the WAT materials. Table 1 illustrates the specific search queries for five selected digital libraries.
Table 1. Selected Digital Libraries, with Search Queries
No. Name Search Query
1 ACM [[Title: ‘‘web’’] OR [Title: ‘‘browser based’’]] AND [[Title: ‘‘test’’] OR [Title: ‘‘testing’’] OR [Title: ‘‘measure’’] OR [Title: ‘‘measurement’’] OR [Title: ‘‘measuring’’] OR [Title: ‘‘check’’] OR [Title: ‘‘checking’’] OR [Title: ‘‘detect’’] OR [Title: ‘‘detecting’’] OR [Title: ‘‘detection’’]]
2 Elsevier ((‘‘web’’) OR (‘‘browser based’’)) AND ((‘‘test’’) OR (‘‘testing’’) OR (‘‘measure’’) OR (‘‘check’’) OR (‘‘checking’’) OR (‘‘detect’’) OR (‘‘detecting’’))
3 IEEE (‘‘Document Title’’:‘‘web’’ OR ‘‘Document Title’’:‘‘browser based’’) AND (‘‘Document Title’’:‘‘test’’ OR ‘‘Document Title’’:‘‘testing’’ OR ‘‘Document Title’’:‘‘measure’’ OR ‘‘Document Title’’:‘‘measurement’’ OR ‘‘Document Title’’:‘‘measuring’’ OR ‘‘Document Title’’:‘‘check’’ OR ‘‘Document Title’’:‘‘checking’’ OR ‘‘Document Title’’:‘‘detect’’ OR ‘‘Document Title’’:‘‘detecting’’ OR ‘‘Document Title’’:‘‘detection’’)
4 Springer ‘‘web’’ AND (‘‘test’’ OR ‘‘testing’’ OR ‘‘measure’’ OR ‘‘measurement’’ OR ‘‘measuring’’ OR ‘‘check’’ OR ‘‘checking’’ OR ‘‘detect’’ OR ‘‘detecting’’ OR ‘‘detection’’)
5 Wiley ‘‘(‘‘web’’) OR (‘‘browser based’’)’’ in Title and ‘‘(‘‘test’’) OR (‘‘testing’’) OR (‘‘measure’’) OR (‘‘measurement’’) OR (‘‘measuring’’) OR (‘‘check’’) OR (‘‘checking’’) OR (‘‘detect’’) OR (‘‘detecting’’) OR (‘‘detection’’)’’ in Title
The survey included studies written in English that were focused on WAT, but excluded academic theses (e.g., master’s and doctoral dissertations), books, posters, and previous survey studies. In addition, some processes relevant to WAT, such as failure diagnosis and regression testing for web applications, were excluded from the survey. The inclusion criteria were:
  • The paper was written in English.
  • The paper was related to WAT.
  • The paper was not a thesis.
  • The paper was not a survey or a systematic literature review.
  • The paper was freely available in open access.
After duplicates were removed and the exclusion criteria applied, the initial set of candidate studies was significantly refined. A snowballing approach (Golmohammadi et al., 2023), which involved examining the references lists of in the selected studies, led to the identification and inclusion of several additional papers. Table 2 summarized the details of the search and filtering process. Eventually, 314 papers were included for the preliminary review and statistical analysis.
Although some papers may have been omitted due to the focus on a subset of reputable publishers, we are confident that the overall trends we report on are accurate and provide a fair representation of the current state of the art in WAT.
3.3. Data Extraction and Collection
Key information was gathered from each study regarding the research motivation, contributions to the field, empirical-evaluation details, common misconceptions, and remaining challenges in WAT. These were reviewed and verified by co-authors of the survey, minimizing the risk of overlooking critical information, and reducing potential errors in the analyses.
We organized the RQ data collection methods as follows:
  • RQ1: Fundamental information for each paper such as publication year, type of paper, and paper source.
  • RQ2: Testing AUT, testing objective, and test case generation techniques.
  • RQ3: Execution methods, tools and platforms, environments, tested items, dataset URL, and open-source status.
  • RQ4: Evaluation metrics, methods and approaches, and tools and techniques.
  • RQ5: Tool name, type, main features and capabilities, URL, and open-source status.
  • RQ6: Overview of the remaining challenges.
Table 2. Selection Results of Primary Studies
No. Name
No. of studies from the search
keywords-based results
No. of studies excluded
based on the selection criteria
No. of studies after
the selection criteria
1 ACM 232 138 94
2 Elsevier 27 19 8
3 IEEE 343 171 172
4 Springer 65 33 32
5 Wiley 19 11 8
Total 686 372 314
4. ANSWER TO RQ1: Publication Evolution and Distribution
This section addresses RQ1 through a detailed analysis of publication trends and their distribution across the literature sources. Our methodology provides a comprehensive analysis of the evolution and distribution of WAT research over the past decade.
4.1. Publication Trends
We collected publication year data from 314 papers (as illustrated in Figure LABEL:Fig:publication_trends) to highlight trends in WAT research from 2014 to 2023.
Figure LABEL:Fig:number_of_publications_per_year shows the annual number of publications, and Figure LABEL:Fig:cumulative_number_of_publications_per_year shows the cumulative total. Figure LABEL:Fig:number_of_publications_per_year shows the fluctuating (but relatively steady) nature of WAT-related publications from 2014 to 2023. The annual publication counts exhibit some variation over the decade, with 32 publications in 2014, followed by slight fluctuations: 35 in 2015, 25 in 2016, 27 in 2017, 29 in 2018, and 34 in 2019. In 2020, the count was 31, which increased to 38 in 2021. This was followed by a decrease to 28 in 2022 and a subsequent rise to 35 in 2023. This pattern reflects a generally stable trend in WAT research output over the years.
The cumulative publication data presented in Figure LABEL:Fig:cumulative_number_of_publications_per_year reflects a consistent growth in research contributions. The trend in Figure LABEL:Fig:cumulative_number_of_publications_per_year suggests that WAT has maintained a steady pace of development, driven by ongoing technological advancements and sustained research interest in the field.
4.2. Paper Source and Publisher
The papers analyzed in this study were from various journals and conferences. As shown in, the majority of the publications were from conferences (71%), with a smaller proportion being from journals (29%). The data presented in Figure LABEL:Fig:venues_distribution_per_year reveals distinct patterns in publication volumes over the years. For instance, conference papers exhibited fluctuations, starting at 25 in 2014, rising slightly to 27 in 2015, and then decreasing to 17 in 2018 before recovering to 26 in 2023. In contrast, journal publications, while also fluctuating, reached their peak of 15 papers in 2019, demonstrating a steadier but less pronounced growth compared to conference papers.
5. ANSWER TO RQ2: Test Case Generation
This section addresses RQ2 by systematically examining key aspects of WAT test case generation, including the Applications Under Test (AUTs) and datasets, testing objectives, and test case generation methods.
5.1. AUTs and Datasets
5.1.1. AUTs
This section reviews the types of AUTs used in WAT, referencing the selected research papers. In WAT, AUTs refer to the web platforms or systems being evaluated (for functionality, performance, security, and usability). The diversity of WAT AUTs emphasizes the need for tailored testing approaches to address specific challenges across different domains: E-commerce platforms, for example, require stringent security measures, whereas government portals may focus on accessibility.
  • E-commerce Platforms: E-commerce applications such as Magento, Shopify, and custom-built online stores are popular WAT AUTs. This may be due to their complex transactional workflows, which involve user authentication, shopping carts, and payment gateways (Pratama and Rhusuli, 2022). Testing often focuses on functional correctness and security, with particular attention to vulnerabilities to potential attacks like SQL Injection (SQLi) and Cross-Site Scripting (XSS) (Bozic et al., 2015a). Performance testing is also essential to ensure that the platforms can handle high traffic volumes during peak times (Wan et al., 2019).
  • Content Management Systems (CMSs): CMS platforms such as WordPress, Drupal, and Joomla are frequently chosen as AUTs because of their modular and extensible nature, which can potentially make them vulnerable, through third-party plugins (Garn et al., 2014; Arcuri et al., 2023). Testing typically emphasizes plugin vulnerabilities, core stability, and validating performance after updates (Li et al., 2014c).
  • Government and Public Service Portals: Government web applications, such as university and public service portals, are often tested for accessibility, usability, and security (Offutt and Thummala, 2019). These platforms often need to comply with accessibility standards such as the Web Content Accessibility Guidelines (WCAG) (Ara et al., 2024), making accessibility testing critical (Vithanage and Jeyamohan, 2016). Security testing emphasizes the prevention of unauthorized access and data breaches, as illustrated by frameworks targeting GDPR compliance in web portals (Amanzholova et al., 2021).
  • Social Media Platforms: Social media platforms, like Facebook and Twitter, pose unique challenges for WAT, due to their real-time interaction and dynamic content. Testing for these applications often examines performance under heavy user loads, cross-browser compatibility, and session management vulnerabilities (Fan et al., 2023; Appelt et al., 2018).
  • Web Application Firewalls (WAFs): WAFs play a key role in defending web applications from attacks like SQLi, XSS, and Cross-Site Request Forgery (CSRF) (Pelivani and Cico, 2021; Zhang et al., 2015a). Testing WAFs involves simulating real-world attack scenarios to evaluate how effectively they distinguish between legitimate and malicious traffic (Zhao and Gong, 2015; Chang et al., 2023).
Table 3. An overview of datasets.
Dataset Name Overview Resource Link Reference
Software Assurance Reference Dataset (SARD) It is a dataset consisting of PHP test cases, where each test case represents a vulnerable or non-vulnerable SQL-related code snippet. The test cases include examples of user input propagating through the application code until a sensitive function (e.g., a database query) is executed. https://samate.nist.gov/SARD/index.php (Fidalgo et al., 2020)
Near-Duplicate Study DataSet It is a dataset of about 100,000 annotated pairs of web pages from open-source web applications and real websites. Each pair is labeled as distinct, near-duplicate, or clone to facilitate the evaluation of near-duplicate detection methods. https://zenodo.org/records/3376730  (Corazza et al., 2021)
Website Screenshots Dataset from Roboflow It is a dataset of 1,206 screenshots from popular websites with 54,215 UI element annotations such as buttons, text, images, links, titles, fields, iframes, and labels. This dataset is used to train object detection models and evaluate the visual consistency of web applications. https://public.roboflow.com/object-detection/website-screenshots  (Aditya et al., 2023)
PhishTank dataset It is a widely used dataset containing known phishing URLs. The dataset provides examples of real phishing attempts that the system uses to test its effectiveness. http://www.phishtank.com  (Nirmal et al., 2021)
Web injection publicly available datasets The datasets consist of various HTTP request samples labeled according to the type of attack they represent, including SQL injection, XSS, command injection, and path traversal. SQL Injection: (1) https://www.kaggle.com/syedsaqlainhussain/sql-injection-dataset,
(2) https://github.com/mhamouei/rat_datasets, XSS Injection: https://www.kaggle.com/syedsaqlainhussain/cross-site-scripting-xss-dataset-for-deep-learning, HTTP Parameters: https://github.com/Morzeux/HttpParamsDataset
 (Román-Gallego et al., 2023; Amouei et al., 2021)
Mitch Framework HTTP Request Dataset It is a dataset based on the usage of the Mitch framework, which contains 5,828 HTTP requests but is reduced to 1901 records after balancing. https://github.com/alviser/mitch  (Beltrano et al., 2023)
NASA 1995 Apache Web Log It is a dataset containing log files from NASA web servers, recording HTTP requests from real users during July 1995. https://ita.ee.lbl.gov/html/contrib/NASA-HTTP.html  (Pavanetto and Brambilla, 2020)
DBpedia It is a structured knowledge base that extracts information from Wikipedia. It contains millions of entities represented as RDF triples. https://www.dbpedia.org/  (Mariani et al., 2014)
fuzzdb It is an open-source database containing a large number of attack patterns. https://github.com/fuzzdb-project/fuzzdb  (Zhang et al., 2017)
ATO-data It is a collection of keystroke dynamics data collected under controlled conditions in a lab and anonymized data from 187,332 sessions of financial web pages, which supports analyzing the changes in different user behaviors and is used to evaluate the effectiveness of imposter detection models. https://github.com/mluckner/ATO-data.git  (Grzenda et al., 2023)
Web Accessibility Evaluation Data It is an automated web accessibility assessment of nearly 3 million web pages obtained between March and September 2021 using Qual Web. https://zenodo.org/records/7494722  (Martins and Duarte, 2023)
PHP-Webshell-Dataset It is a dataset containing 2,917 samples from 17 webshell collection projects. https://github.com/Cyc1e183/PHP-Webshell-Dataset  (Gogoi et al., 2022)
PHP Vulnerability Test Suite It is a dataset containing 9408 PHP source code samples, including 5600 safe samples and 3808 unsafe samples, which is widely used for the detection of XSS vulnerabilities and the training of prediction models. https://github.com/stivalet/PHP-Vulnerability-test-suite  (Gupta et al., 2015)
Datasets of three common web attack payloads (SQLi, XSS, RCE) It is a dataset generated by attack grammar, including SQL injection, cross-site scripting, and remote command execution attack payloads, used to train and test the performance of GPTFuzzer on WAF. https://github.com/hongliangliang/gptfuzzer  (Liang et al., 2023)
Joomla! and MantisBT test suite It is a data set that contains multiple representative use cases, covering functions such as article management and user management. It has 47 manually written test cases and 453 test steps. https://zenodo.org/record/4973219  (Kirinuki et al., 2021)
GHTraffic It is a dataset of significant size comprising HTTP transactions extracted from GitHub data (i.e., from 04 August 2015 GHTorrent issues snapshot) and augmented with synthetic transaction data. https://zenodo.org/record/1034573  (Bhagya et al., 2019)
5.1.2. Datasets
Datasets are essential in WAT, providing the inputs needed to simulate real-world usage and evaluate application behavior across key areas such as security, functionality, and performance. The datasets are often comprised of user data, predefined test cases, or automatically generated data. They support a range of testing scenarios, including edge cases, aiming to identify potential vulnerabilities and performance bottlenecks.
Each WAT dataset was designed to meet specific testing objectives, allowing for targeted assessments of various aspects. For example, the Software Assurance Reference Dataset (SARD) (Fidalgo et al., 2020) can help to identify security vulnerabilities, such as SQLi, thus contributing to more secure software development practices. The PhishTank dataset (Nirmal et al., 2021), which contains real-world phishing URLs, can support evaluation of phishing detection systems, enhancing web application security.
Datasets like the Website Screenshots Dataset (Aditya et al., 2023) facilitate the assessment of visual consistency and UI accuracy, helping to ensure an optimal user experience. The NASA web log data (Pavanetto and Brambilla, 2020) provides real-world HTTP requests, offering valuable insights into traffic patterns and server performance under various loads. DBpedia (Mariani et al., 2014) supports the validation of semantic web applications: It provides structured data that can be used to help verify data integration and usage.
These datasets enable testers to adopt focused testing strategies that ensure a thorough evaluation of web applications across multiple dimensions. Table 3 summarizes key datasets used in WAT, each with a distinct focus, ranging from security testing.
5.2. Test Objectives
WAT test objectives include detecting bugs, ensuring security, evaluating performance, assessing the functionality of Graphical User Interfaces (GUIs), and identifying vulnerabilities (such as for XSS and SQLi).
Figure 3. Proportions of test objectives in surveyed papers.
Figure 3 shows the proportion of the different objectives across the surveyed literature. Web vulnerability detection, including XSS and SQL Injection detection, accounted for the largest proportion (29.94%), emphasizing the necessity of addressing critical security risks. Security testing represented 27.71%, highlighting its importance in protecting applications from security threats. GUI testing accounted for 19.43%. Performance evaluation constituted 12.42%, and bug/defect detection represented 10.51%. The following sections provide a detailed analysis of each testing objective.
5.2.1. Bug/Defect Detection
Detecting and correcting defects is essential for maintaining web application integrity. Techniques such as combining unsupervised learning with embedding layers effectively optimize test case prioritization while reducing redundancy (Demirel and Sandikkaya, 2023). Modern web applications often involve intricate interdependencies between components, requiring advanced techniques to detect defects across diverse states and modules. This complexity necessitates strategies such as dynamic test case generation and static analysis to ensure robust coverage (Maskur and Asnar, 2019).
Static code analysis can complement regression testing, and may identify issues early in development. Tools that use static analysis (such as taint analysis) detect potential errors in the source code without execution (Maskur and Asnar, 2019; Nunes et al., 2019). For example, the Q-learning-based Selenium-Java tool dynamically explores web GUI elements and learns optimal test paths, significantly improving testing coverage for complex systems (Fan et al., 2023).
5.2.2. Security Testing
The goal of security testing is to ensure that web applications remain protected against unauthorized access, data breaches, and other security threats. Security testing provides a systematic framework for evaluating the robustness of applications through the identification and mitigation of vulnerabilities before and during runtime (Medeiros et al., 2016).
To achieve comprehensive protection, security testing employs a combination of methodologies. Static analysis examines source code for vulnerabilities, such as unhandled exceptions and unsanitized inputs, enabling early detection of issues during development (Zhao and Gong, 2015). Dynamic analysis evaluates applications during runtime, simulating potential attack scenarios to uncover vulnerabilities that static analysis might overlook (Pellegrino et al., 2015). Hybrid analysis combines static and dynamic methods to address their individual limitations, offering enhanced coverage and reducing false positives (Guan et al., 2021). Automated tools, such as Burp Suite, enhance the efficiency of dynamic analysis by enabling real-time evaluation of application defenses (Garn et al., 2014).
5.2.3. Performance Evaluation
Performance evaluation aims to ensure that an application remains efficient and stable under varying load conditions (Pelivani and Cico, 2021). Essential to this are load testing and stress testing, which simulate diverse user traffic scenarios to help developers identify performance bottlenecks, especially in environments with high concurrency (Pelivani and Cico, 2021; Lounis et al., 2014).
Advanced tools such as JMeter (Kiran et al., [n. d.]; Tiwari et al., 2023) and LoadRunner (Cai and Hu, 2014) can automate measurement of some important metrics, including response times, throughput, and system resource utilization (e.g., CPU, memory, disk I/O) (Fan et al., 2023). They can do this across a spectrum of load levels (Fan et al., 2023). These tools can provide comprehensive insights into system behavior under stress, helping the optimization of the system architecture, and ensuring reliability during peak load conditions (Leotta et al., 2020).
5.2.4. GUI Testing
The aim of GUI testing is to ensure that the application’s user interface remains consistent, in terms of both functionality and user experience, across different devices and browser environments (Zhang et al., 2015b). Automated tools can simulate user interactions, exploring the GUI’s responsiveness under diverse conditions (Lin et al., 2017): Cross-browser testing with tools like Selenium Grid (Kiranagi and Shyam, 2017), for example, can help to ensure compatibility across different platforms (Pelivani and Cico, 2021). Meanwhile, Scout uses gamified features (Fulcini and Ardito, 2022), including progress tracking and dynamic error injection, to encourage systematic exploration and improve test coverage.
Image comparison and computer vision technologies are also extensively used to detect interface changes, and to identify potential UI errors (Leotta et al., 2020). These tools capture and compare screenshots from different stages of testing, ensuring that updates do not introduce unintended changes to the user interface (Zhao and Gong, 2015). The recent integration of AI in testing can further enhance efficiency by learning from user interactions and predicting areas that may be likely to fail (Fan et al., 2023; Chang et al., 2023).
5.2.5. Web Vulnerability Detection
Web vulnerability detection focuses on identifying and mitigating specific threats in web applications, such as Cross-Site Request Forgery (CSRF), directory traversal, Cross-Site Scripting (XSS), and SQL Injection (SQLi). These vulnerabilities, if left unaddressed, can compromise application functionality, data security, and user privacy (Ben Jaballah and Kheir, 2016; Zhang et al., 2015a).
To detect these vulnerabilities, static analysis examines source code to identify input validation flaws and hardcoded credentials, enabling early detection of coding issues (Zhao and Gong, 2015; Lounis et al., 2014). Dynamic analysis simulates attack scenarios during runtime, evaluating the application’s behavior under unexpected interactions (Ben Jaballah and Kheir, 2016). Hybrid analysis combines the strengths of static and dynamic approaches, providing enhanced detection accuracy and reducing false positives (Guan et al., 2021).
Modern advancements in vulnerability detection integrate machine learning and deep learning techniques. Methods based on Recurrent Neural Networks (RNNs) analyze input and output patterns to detect malicious behaviors, demonstrating high effectiveness in identifying XSS and SQLi attacks (Guan et al., 2021). Additionally, OWASP ZAP is widely used to automate the identification of vulnerabilities, offering strong support for detecting SQLi and XSS (Zhang et al., 2015a).
5.3. Generation Methods
WAT can make use of many test case generation methods, including: Model-Based (Dias Neto et al., 2007), Combinatorial and Optimization-Based (Nie and Leung, 2011), Learning-Based (Alagarsamy et al., 2024), Data and Content-Driven (Elbaum et al., 2008), Crawling and Exploration-Based (Mesbah et al., 2012), Template and Page Object-Based (Stocco et al., 2016), and Security Test Case Generation techniques (Felderer et al., 2016). These methods offer a broad selection of strategies that can address diverse testing needs.
Figure 4. Proportions of generation methods in surveyed papers.
Figure 4 presents the proportions of some key WAT test generation methods in the surveyed literature. Model-Based Testing (MBT), which uses structural models (e.g., UML, statecharts) that reflect system behavior, is the most prominent, appearing in 27.50% of the papers. Combinatorial and Optimization-Based Testing (20.00%) enhances coverage through input-variable combinations. Data and Content-Driven Testing (20.00%) make use of input validation across diverse scenarios. Security Testing was found in 15.00% of the papers. Learning-Based Testing (10.00%), Template and Page Object-Based Testing (5.00%), and Crawling and Exploration-Based Testing (2.50%) were less commonly reported.
5.3.1. MBT
Model-Based Testing (MBT) automatically generates test cases, enhancing both efficiency and coverage. UML diagrams, statecharts, ER diagrams, and BPMN models are commonly used to represent the system, and to generate the concrete test cases. Panthi & Mohapatra (Panthi and Mohapatra, 2017) and Suhag & Bhatia (Nabuco and Paiva, 2014) used UML models; Indumathi & Begum (Indumathi and Begum, 2016) used ER diagrams and state transition models for database testing. Akpinar et al. (Akpinar et al., 2020) showed how MBT could enhance coverage of complex interactions by integrating GraphWalker, Selenium, and JUnit. Moura et al. (de Moura et al., 2017) used BPMN models to generate automated test scripts that covered functional test paths in Business Process Management (BPM) applications. Wang et al. (Wang et al., 2016b) proposed an approach based on UML statecharts for generating Web link security test scenarios. Thummala & Offutt (Thummala and Offutt, 2016) used Petri nets to detect concurrency behaviors. Fard et al. (Milani Fard et al., 2014) and Waheed et al. (Waheed et al., 2020) showed how model inference and script automation could enhance the automation and test coverage.
5.3.2. Combinatorial and Optimization-Based Testing
Techniques based on combinatorial or optimization can generate efficient test cases, which can significantly reduce the number of tests while improving coverage. Qi et al. (Qi et al., 2017) introduced an automated approach based on combinatorial strategies: This enhanced the coverage of dynamic pages. Bozic et al. (Bozic et al., 2015b) reported on the effectiveness of IPOG and IPOG-F combinatorial testing algorithms for detecting XSS vulnerabilities. Wang et al. (Wang et al., 2023) and Wang et al. (Wang et al., 2019) applied evolutionary algorithms and combinatorial testing methods to generate test cases covering complex paths in Web applications: Their work demonstrated the potential for genetic algorithms and simulated annealing in Web testing. Wang et al. (Wang et al., 2016a) used combinatorial testing to construct navigation maps for dynamic Web applications, enhancing the test coverage and effectiveness.
5.3.3. AI-based Testing
Testing methods can also use machine learning and artificial intelligence (AI) techniques to analyze and generate Web application behaviors and inputs, automating the creation of effective test cases. Chang et al. (Chang et al., 2023) proposed the WebQT tool based on reinforcement learning, significantly improving testing efficiency and coverage through intelligent path selection. Pavanetto & Brambilla (Pavanetto and Brambilla, 2020) used Generative Adversarial Networks (GANs) to generate highly realistic Web navigation paths, effectively simulating real user operations. Chaleshtari et al. (Chaleshtari et al., 2023) demonstrated the potential of AI techniques in security testing by using mutation testing to detect security vulnerabilities in Web systems. Shahbaz et al. (Shahbaz et al., 2015) introduced a technique for generating test data based on regular expressions and Web searches, producing valid and invalid WAT data. Li et al. (Li et al., 2014a) developed the SymJS framework, which combines symbolic execution and dynamic feedback to improve the coverage and accuracy of JavaScript WAT.
5.3.4. Data and Content-Driven Testing
Data and content-driven testing methods aim to cover all possible input combinations, to comprehensively verify application functionality and security. Hanna & Jaber (Hanna and Jaber, 2019) proposed a method based on analyzing client-side user input fields, effectively covering various input scenarios. Hanna & Munro (Hanna and Munro, 2018) used semantically invalid input data to ensure that Web applications reject inputs that do not meet semantic constraints, thereby enhancing security. Shahbaz et al. (Shahbaz et al., 2015) combined regular expressions with Web searches to automatically generate test cases, verifying string validation logic in Web applications. Nagowah & Kora-Ramiah (Nagowah and Kora-Ramiah, 2017) developed the CRaTCP tool, which automatically extracts control combinations and generates test cases, achieving comprehensive WAT coverage. Mariani et al. (Mariani et al., 2014) utilized the data from applications to automatically generate valid input data (which satisfies the syntactical and semantical requirements) to test complex web applications and ensure the coverage of input combinations.
5.3.5. Crawling and Exploration-Based Testing
Methods based on crawling and exploration are particularly suited for testing the complex navigation and state changes of dynamic Web applications. Fard et al. (Milani Fard et al., 2014) combined automated crawling with manually written test cases to enhance coverage and explore unexplored paths. Santiago Balera and Junior (Balera and de Santiago Júnior, 2022) proposed a multi-perspective crawling approach using generative hyper-heuristic algorithms to capture diverse user behaviors and generate comprehensive test cases. Yousaf et al. (Yousaf et al., 2019) integrated model-driven techniques with crawling strategies to navigate interfaces and improve test case quality systematically. Jesus et al. (de Jesus et al., 2015) leveraged task-driven methods to generate navigation paths aligned with functional workflows. Nguyen et al. (Nguyen et al., 2022) applied Q-learning to explore dynamic state spaces, effectively detecting vulnerabilities like XSS. Pavanetto and Brambilla (Pavanetto and Brambilla, 2020) utilized deep learning models, combining RNNs and GANs to simulate realistic navigation paths, enhancing test coverage and interaction fidelity.
5.3.6. Template and Page Object-Based Testing
Robust and efficient test cases can be generated using templates or the Page Object Model. Yu et al. (Yu et al., 2015) used an incremental testing approach based on the Page Object Model that automatically generates page objects to represent Web pages. Neves et al. (de Almeida Neves et al., 2022) developed the Morpheus Web Testing tool, which extracts UI component information from JSF and Primefaces frameworks to generate system-level functional test cases, effectively covering complex user interaction scenarios. Leotta et al. (Leotta et al., 2015) utilized meta-heuristic techniques, including greedy and genetic algorithms, to automatically generate robust XPath locators, which reduces the maintenance overhead caused by DOM evolution.
5.3.7. Security Testing Generation Techniques
There are also techniques for generating test cases to detect security vulnerabilities in Web applications. Common vulnerabilities include for XSS and SQLi attacks. Akrout et al. (Akrout et al., 2014) proposed an automated black-box Web vulnerability identification and attack scenario generation method, enhancing vulnerability detection and accuracy. Awang et al. (Awang et al., 2019) used Cartesian product algorithms (Agesen, 1995) to generate SQLi test data, effectively detecting vulnerabilities. Wang et al. (Wang et al., 2016b) used UML statecharts to generate Web link security test scenarios, enhancing robustness in complex security environments. Nguyen et al. (Nguyen et al., 2022) used Q-learning algorithms (Jang et al., 2019), detecting security vulnerabilities, and demonstrating the potential for intelligent algorithms in Web security testing. Wang et al. (Wang et al., 2019) used evolutionary testing, combining client-side and server-side testing scenarios to coverage and efficiency, and demonstrating the potential for combinatorial optimization algorithms in Web security testing.
6. ANSWER TO RQ3: Test Case Execution
This section addresses RQ3, examining the execution of test cases within the context of WAT. The section explores the various methods, tools, platforms, environments, and key considerations of WAT testing: Each aspect is analyzed with a focus on its practical application, and its contribution to test-execution effectiveness.
6.1. Environments
The selection of testing environments is critical for WAT effectiveness, directly influencing the accuracy and scope of performance assessments. According to the literature, testing environments can be categorized into local (Wang et al., 2023), remote (Calzavara et al., 2021), and cloud-based (Wei et al., 2019).
Figure 5. Proportions of execution environments in surveyed papers.
Figure 5 shows that local environments were the most frequently used, accounting for 87.58%. The control and immediate feedback potential of local environments make them particularly suitable for iterative testing in early development (Everett and McLeod Jr, 2007). Remote environments (4.78%) support testing under various network conditions: This is essential for applications with geographically dispersed users. Cloud-based environments (7.64%) offer scalability and flexibility to meet complex testing demands and support continuous integration in large-scale applications.
6.1.1. Local Environments
The initial stages of development and testing often take place in local environments, which provide a controlled and isolated setting that facilitates rapid iteration and debugging. In these environments, because the web applications are deployed on the same machine that the tests are being executed on, the speed of iteration and bug fixing can be increased. Tools like Apache JMeter are increasingly adopted for testing web applications in local environments (Tiwari et al., 2023). JMeter excels in performance and functional testing and provides a reliable framework to validate isolated components during unit and integration testing, especially in controlled local settings.
A challenge for local environments, however, is that they cannot easily some simulate real-world scenarios like network latencies, distributed system behaviors, and large-scale performance (Prisadi et al., 2023). To address these limitations, tools like Docker have been widely used to create containerized environments that maintain consistency across different stages of development and production (Halin et al., 2019). Docker ensures that testing environments are uniform, thus reducing discrepancies among development, testing, and production setups (Marquardt and Buhl, 2021).
6.1.2. Remote Environments
Remote environments extend local capabilities by enabling the execution of test cases across multiple physical or virtual machines. Tools like Selenium Grid (Balera and de Santiago Júnior, 2022) play an important role in remote environments, enabling cross-browser and cross-platform testing (Debroy et al., 2018). Selenium Grid helps ensure that web applications function consistently across different environments and configurations (Jana and Oprea, 2019).
Despite their advantages, remote environments also face challenges, including network latency and the need to synchronize test execution across different systems (Zhou et al., 2014). Docker, in conjunction with Selenium Grid, helps alleviate these issues by ensuring that the test environments are consistent across remote nodes (Halin et al., 2019).
6.1.3. Cloud-Based Environments
Cloud-based environments offer the most flexible and scalable solution for large-scale WAT. Platforms such as Amazon Web Services (AWS), Google Cloud Platform (GCP), and Microsoft Azure offer dynamic resource allocation, allowing tests to be executed on-demand across multiple geographical locations (Ludinard et al., 2018; Su and Li, 2021). A cloud environment can simulate real-world scenarios (including high traffic and various network conditions), and can support global users (Arcuri et al., 2023). AWS Lambda, for example, is an important tool for cloud environments, providing a serverless architecture that allows testing to run without manual infrastructure management (Mukherjee et al., 2014). It can also dynamically expand resources based on test execution requirements, ensuring efficient test runs.
6.2. Execution Methods
Test-case execution has a critical role in the verification of web application functionality, security, and performance. Various methods can be used to execute test cases, each with certain advantages, depending on the nature and complexity of the application. This section provides a detailed exploration of the three primary approaches to test-case execution: manual, automated, and hybrid. Each approach is analyzed for its practical application, highlighting any contributions to the overall effectiveness of WAT test execution.
Figure 6. Proportions of execution methods in surveyed papers.
Figure 6 presents an overview of the s of different approaches to test-case execution in the surveyed literature. Automated execution dominates, appearing in 91.08% of the literature. This is as expected, given its efficiency and scalability for repetitive tasks and regression testing. Automated execution can support tasks requiring consistency across environments. Manual execution, although only in 1.91% of the surveyed papers, remains essential for tasks like exploratory testing and scenarios where human intuition is required. Hybrid execution, accounting for 7.01%, combines both approaches, optimizing test coverage and flexibility by automating routine tasks while leveraging manual testing for subjective analysis. This can be particularly valuable in large-scale applications.
6.2.1. Manual Execution
Manual execution involves human testers directly interacting with web applications to verify their behavior. This can be particularly effective for exploratory testing, as testers need flexibility to dynamically respond to application behavior and detect unexpected issues during testing. Exploratory testing can help to identify user-experience issues, or unexpected errors that are difficult to predict when designing automated scripts (Fan et al., 2023).
Manual testing is also often used to evaluate new features or complex user interfaces, which can require human judgment to assess usability, aesthetics, and overall user satisfaction (Leotta et al., 2020). However, there can be significant limitations to manual testing, including that it can be time-consuming, labor-intensive, and prone to human errors. This can be especially the case when dealing with large applications that require the execution of hundreds of test cases (Garg et al., 2015). Therefore, although manual testing remains essential for specific tasks, it may be most suitable for early development stages or areas that require subjective validation (Gupta and Bajpai, 2014).
6.2.2. Automated Execution
Automated execution is a more scalable and efficient approach, especially when repeating test cases in different environments or during regression testing. Typically, scripts and tools are used to execute test cases, capture results, and compare them with predefined expected outcomes (Leotta et al., 2020). This is often done without need for further human intervention. Selenium, for example, is widely used to automate functional testing across various browsers, ensuring consistent behavior and compatibility (Appelt et al., 2018).
Automated execution can significantly improve efficiency by reducing human involvement and allowing tests to run continuously in different environments. However, automated execution can require considerable upfront investment in scripting and tool configuration (Ammann and Offutt, 2017). Furthermore, it may not be suitable when subjective analysis is required, such as for visual-interface testing (Garg et al., 2015; Lounis et al., 2014).
6.2.3. Hybrid Execution
Hybrid execution combines both manual and automated techniques, drawing on the strengths of both approaches. In hybrid models, manual testing is typically used for exploratory tasks or validation of new features, while automated scripts handle repetitive tasks and regression testing  (Pelivani and Cico, 2021). This approach offers flexibility by allowing testers to manually address complex, subjective scenarios while taking advantage of automation for the repetitive, time-consuming processes (Guan et al., 2021; Garg et al., 2015).
It has been reported that combining UML activity diagrams with manual and automated test cases could improve overall test coverage and shorten execution time (Zhao and Gong, 2015). Similarly, a hybrid approach can be adopted in security testing, where potential vulnerabilities (such as SQLi points) can be detected automatically, and manual testers can then validate the findings (Singh et al., 2020). The combination of the thoroughness of manual testing and the efficiency of automation makes hybrid execution often the most practical solution for complex and large-scale web applications (Gupta and Bajpai, 2014).
7. ANSWER TO RQ4: Evaluations and Metrics
This section addresses RQ4 by exploring WAT metrics and evaluation methods. We first discuss the primary effectiveness metrics, such as test case quality, coverage, failure diagnosis, and vulnerability detection. This is followed by an analysis of efficiency metrics, such the time required for generation, execution, and error detection. Finally, empirical studies are reviewed to explain how these metrics have been applied to evaluate and improve WAT effectiveness and efficiency.
7.1. Test Effectiveness
Test effectiveness evaluates the ability of WAT methodologies to identify faults, vulnerabilities, and failures. This involves the use of specific metrics, including test-case quality, code coverage, failure-detection capability, and vulnerability-detection effectiveness. These metrics provide insights into the thoroughness and accuracy of WAT testing processes.
Figure 7. Proportions of test effectiveness metrics in surveyed papers.
Figure 7 presents an overview of the different test effectiveness metrics used in the surveyed literature. Failure-detection capability is the most frequently used metric, accounting for 30.22% of the papers. Code and coverage metrics (25.75%) measure the breadth of testing. Vulnerability detection comprises 20.90%, while test-case quality accounts for 16.42%. A small proportion of the literature, 6.72%, involved other metrics: For example, Kumar et al. (Saharan et al., 2018), and Tiwari et al. (Tiwari et al., 2023), for example, refer to performance-based metrics such as response time, error rate, and throughput.
7.1.1. Test-Case Quality
The quality of test cases directly affects the ability to detect defects in web applications. Dynamic testing approaches, such as Q-learning-based frameworks, have demonstrated substantial improvements in test-case adaptability. These methods optimize element interaction paths, increase branch and element coverage, and effectively identify critical defects in complex application environments (Fan et al., 2023). These tools provide a scalable solution for testing modern web applications by enhancing exploratory capabilities.
Prioritizing test cases based on risk and modification frequency has enhanced defect detection efficiency. These strategies uncover critical defects earlier by focusing on high-risk and frequently modified components, which streamlines testing processes and improves efficiency and effectiveness (Ali-Shahid and Sulaiman, 2015).
Furthermore, fast evaluation techniques that leverage clustering-based sampling can improve test-case efficiency in large-scale applications. These methods group similar application paths and ensure representative coverage, which significantly reduces redundant testing efforts while maintaining reliability (Zhang et al., 2015b).
7.1.2. Code Coverage Metrics
Code coverage metrics — including line, branch, and function coverage — can help show how thoroughly an application’s codebase is exercised during testing (Nguyen et al., 2019; Wang et al., 2016a). Coverage analysis tools like Selenium allow testers to capture interactions across multiple layers of the web application, from the front-end UI to the back-end logic (Paul and Tommy, 2018).
However, high coverage alone does not guarantee comprehensive fault detection. Research has shown that combining coverage metrics with fault injection methods enhances the ability to uncover defects (Chaleshtari et al., 2023). Injecting controlled faults into the system, and observing whether or not the test cases can detect them, provides a more rigorous assessment of a test suite’s fault-detection capabilities (Gupta and Bajpai, 2014). Mutation testing complements code coverage metrics by introducing controlled changes (i.e., “mutants”) to evaluate the fault-detection capabilities of test cases (Habibi and Mirian-Hosseinabadi, 2015; Leotta et al., 2024). While coverage metrics measure the breadth of testing, mutation testing ensures that covered paths are rigorously tested, enhancing the reliability of the testing process.
7.1.3. Failure-Detection Capability
Failure-detection capability refers to the accurate identification and localization of faults during testing. Techniques such as dynamic trace-based analysis compare the execution traces of different versions of the application to identify discrepancies (Touseef et al., 2019). This approach has proven effective in failure detection, particularly when combined with automated tools for analyzing execution logs and detecting potential issues, such as security vulnerabilities, in large-scale applications. This can help testers to pinpoint the specific lines of code or configuration changes responsible for faults, significantly reducing the time required for debugging and resolution (Walsh et al., 2015; Halin et al., 2019).
Automated diagnostic tools further improve this process by analyzing execution logs and error traces to identify the exact points of failure (Sunman et al., 2022). These tools are particularly useful for large-scale web applications, where manual fault-localization could be time-consuming and error-prone (Fan et al., 2023).
7.1.4. Vulnerability Detection
Vulnerability detection includes the identification of potential security issues such as SQLi and XSS that could be exploited by attackers (Awang et al., 2019; Mokbal et al., 2019). Automated vulnerability detection tools simulate real-world attack scenarios, testing the robustness of the application’s security defenses (Nagpure and Kurkure, 2017).
Recent advances in machine learning have enhanced vulnerability-detection capabilities. These systems analyze traffic and behavioral patterns to identify anomalies that may indicate vulnerabilities (Jana and Oprea, 2019). As machine learning models learn from new data, they continuously improve their ability to detect both known and unknown threats, providing a dynamic defense against emerging security risks (Calzavara et al., 2020).
7.2. Test Efficiency
Test efficiency is a core WAT metric, directly affecting the cost effectiveness of the testing process. Efficiency metrics assess the resources required for comprehensive test coverage and fast fault detection. Some of the main metrics include the time taken for test case generation (Gupta and Bajpai, 2014) and execution (Pelivani and Cico, 2021), and error detection (Fan et al., 2023).
Figure 8. Proportions of test efficiency metrics in survey papers.
Figure 8 presents an overview of the different test efficiency metrics used in the surveyed literature. Test execution time and test case generation time are the most frequently measured metrics, each accounting for 30.83% of the literature. Error detection time is slightly more widely used, appearing in 32.33% of the surveyed papers. A smaller proportion, 6.02%, includes other metrics: For instance, Sung et al. (Sung et al., 2016) used static analysis methods to reduce redundant tests, thereby improving testing speed and efficiency.
7.2.1. Test Case Generation Time
The test-case generation time measures how efficiently test cases are created and is a major part of the overall testing efficiency. Several techniques have been developed to streamline this process. The Page Object Pattern (POP) method has also been able to reduce the test-case generation time as the test suite grows due to the reuse of defined objects. This approach improves efficiency as the same components are used across multiple test cases (Leotta et al., 2020). Studies involving WebQT, which employs machine learning algorithms, have shown that test-case generation can be significantly accelerated through optimized exploration and test-sequence generation (Gupta and Bajpai, 2014). Tools like CRaTCP can automatically generate and execute all possible test case combinations to further reduce the manual efforts (Nagowah and Kora-Ramiah, 2017).
7.2.2. Test-Execution Time
The test-execution time refers to the time taken to execute the generated test cases. Automated testing frameworks like Selenium and Katalon Studio can dramatically reduce execution time (compared to manual testing) (Pelivani and Cico, 2021). Research has confirmed their effectiveness in optimizing test execution times, particularly in large-scale web applications (Pelivani and Cico, 2021). Genetic algorithms, for example, have been used to reduce execution time by optimizing test paths, leading to earlier detection of faults (Stallenberg and Panichella, 2019). Concurrent execution techniques, where multiple test cases are run in parallel, can also further reduce execution times (Garg et al., 2015). Prioritization techniques that execute high-risk test cases first can lead to more efficient use of resources by focusing on the most critical components.
7.2.3. Error-Detection Time
The error-detection time reflects how quickly errors or vulnerabilities are identified during testing. Hybrid methods that integrate static and dynamic analysis improve detection efficiency by focusing on critical code paths and minimizing redundant analysis (Zhao and Gong, 2015). Techniques leveraging anomaly detection through multi-model approaches accelerate fault detection by identifying deviations in real-time and ensuring low false positive rates (Zhang et al., 2017). By minimizing manual intervention, automated test suites enable real-time fault identification and significantly accelerate the detection process (Paul and Tommy, 2018). Moreover, combining diverse static analysis tools has significantly improved detection speed and accuracy, particularly in scenarios with limited resources or time constraints (Nunes et al., 2019).
7.3. Empirical Studies
This section examines the use of empirical studies in WAT, focusing on effectiveness and efficiency metrics. It explains how these studies can enhance testing performance, optimize testing processes, and offer insights for further improvements in testing tools.
7.3.1. Application of Effectiveness and Efficiency Metrics in WAT
Coverage has been widely used to measure testing effectiveness in empirical WAT studies (Garousi et al., 2021; Mattiello and Endo, 2022). MBT can generate high-quality test cases, with good code coverage and fault-detection capabilities (Nabuco and Paiva, 2014; Nguyen et al., 2019). Compared to manual testing, MBT provides a more systematic approach to cover application requirements, especially for complex systems, and achieves better testing efficiency (Wang et al., 2019). MBT is particularly effective at handling the complex interactions between multiple modules or interfaces within large-scale web applications. Empirical studies have shown that MBT reduces test redundancy and optimizes test case generation (Hanna and Munro, 2018).
Automated tools like Model-Based Testing Leveraged for Web Tests (MoLeWe) (Mattiello and Endo, 2022) have demonstrated excellent testing efficiency (Li and Zeng, 2014): MoLeWe can generate high-quality test cases, increasing test coverage while reducing execution time (Garousi et al., 2021). It can significantly reduce manual testing time, especially in applications with frequent version updates. MoLeWe has also demonstrated flexibility in resource allocation, optimizing system resource utilization in multi-threaded testing.
Fault Detection Density (FDD) and Fault Detection Effectiveness (FDE) are core effectiveness metrics used to evaluate the fault-detection capabilities of testing tools (Habibi and Mirian-Hosseinabadi, 2015). The MUTANDIS mutation-testing tool can generate non-equivalent mutants, significantly improving the fault-detection rate in JavaScript applications (Mirshokraie et al., 2014). Mutation testing can covers a wide range of fault scenarios using diverse mutants: This can improve the comprehensiveness of the test cases. MUTANDIS has also been shown to enhance fault-detection efficiency while reducing resource waste caused by redundant testing paths.
Efficiency metrics are not only used for coverage and fault-detection capabilities, but also in execution time and resource utilization (Putri et al., 2017; Ali and Badr, 2015). Empirical studies have shown that tools like JMeter can optimize performance testing through distributed architectures (Saharan et al., 2018; Kiran et al., [n. d.]). By parallelizing the execution of test cases, JMeter can significantly reduce execution time while ensuring system stability under high-load conditions. This architecture has been shown to perform well in real-world scenarios, especially in large-scale multi-user concurrent environments (Guarnieri et al., 2017).
7.3.2. Improvements in Testing Performance through Empirical Studies
Empirical studies have confirmed the effectiveness of reinforcement learning techniques in the prioritization of test cases (Chang et al., 2023): High-risk areas can be prioritized, thereby reducing overall testing time and improving fault-detection efficiency.
Automated tools can also improve testing efficiency. MoLeWe can automatically generate test cases and scripts, minimizing manual intervention and significantly enhancing automation (Mattiello and Endo, 2022): This not only reduces the error rate from manual operations but also enables the rapid adaptation of test suites to accommodate new features or updates.
8. ANSWER TO RQ5: Available Tools
This section addresses RQ5, identifying and cataloging the WAT tools that are currently available. This should provide an overview of the technological resources available to support this field. We first list the tools, and then classify them based on their types and features. The section concludes with some details on tool accessibility, including links for further exploration.
8.1. Testing Tools
Various WAT tools are available to streamline and automate the testing process, supporting efficiency, scalability, and consistency across diverse environments. These tools can support the automation of various different types of testing (including functional, regression, security, and performance testing), helping to examine the reliability and robustness of web applications.
Figure 9 presents an overview of the proportions of testing tools found in the surveyed literature. Security testing tools account for 41.57%, highlighting the increasing focus on mitigating cybersecurity risks. Functional and regression testing tools represent 38.95%, emphasizing their critical role in validating application functionality and maintaining system stability following changes. Performance testing tools comprise 11.34%, and cross-platform/distributed environment tools make up 8.14%, reflecting their specialized contributions to optimizing system performance and ensuring consistency across platforms. This distribution reveals varying priorities in testing, with security and functional testing receiving the most attention, and performance and cross-platform testing appearing to be applied only as required.
Figure 9. Proportions of testing tools in surveyed papers.
8.1.1. Functional and Regression Testing Tools
Functional and regression testing tools, like Selenium (Pelivani and Cico, 2021) and Katalon Studio (Pelivani and Cico, 2021), have been widely used for cross-browser and cross-platform testing: By automating test execution across various browsers and operating systems, they help to verify that web applications function correctly after updates and code changes
8.1.2. Security Testing Tools
Tools like Burp Suite (Nagpure and Kurkure, 2017) and OWASP ZAP (Amankwah et al., 2020) can help to identify security vulnerabilities, including SQLi, XSS, and other common web application threats. These tools are essential for protecting web applications from attacks, and ensuring compliance with security standards.
8.1.3. Performance Testing Tools
JMeter (Fan et al., 2023) and LoadRunner (Khan and Amjad, 2016) have been used to simulate high-traffic loads, helping to evaluate how web applications perform under stress. They assess critical aspects of performance, including response times and throughput, evaluating whether or not the application remains stable under heavy load conditions.
8.1.4. Cross-platform and Distributed Environment Tools
Tools such as Docker (Marquardt and Buhl, 2021) and Selenium Grid (Kiranagi and Shyam, 2017) enable test cases to be executed in distributed environments and containerized platforms. They help to ensure consistency in execution and provide scalability for larger systems, supporting the efficient management of complex testing scenarios.
8.2. Tool Accessibility
The ease of access to WAT tools is a crucial factor in their adoption, and eventual effectiveness in the software development and testing community. Table LABEL:TAB:wat_tools lists some popular WAT tools [^1] To measure the popularity of each WAT tool, we collected the number of times to be mentioned in the selected 314 primary studies., categorized by type, core functionalities, and access links: This structured presentation facilitates the selection of tools based on specific testing requirements, ranging from performance evaluation and security vulnerability detection to automated UI testing and code-quality analysis. Due to page limit constraints, we provide a comprehensive summary of nearly all relevant tools, available in the online supplementary material (Li et al., 2024b).
Table 4. A comprehensive overview of tools.
Tool Name Tool Type Overview Resource Link Reference
Selenium Open-source It is open source, cross-platform, and supports multiple languages and browsers. https://www.selenium.dev/
https://github.com/SeleniumHQ
 (Aditya et al., 2023)
ZAP Open-source It is an open-source web application scanner that excels at finding common web application vulnerabilities. https://www.zaproxy.org/
https://github.com/zaproxy
 (Amankwah et al., 2020)
JMeter Open-source It is an open-source tool for load testing and performance measurement, especially for web applications. It allows the simulation of a large number of users and provides detailed reports. https://jmeter.apache.org/
https://github.com/apache/jmeter.
 (Cai and Hu, 2014)
Arachni Open-source It is a web vulnerability scanner designed to automatically detect and assess security vulnerabilities in web applications. https://github.com/Arachni/arachni  (Román Muñoz et al., 2018)
Crawljax Open-source It is a tool for automatically crawling and testing web applications, which automatically explores the GUI of a web application, creates state flow diagrams, and generates test cases based on interactions. https://github.com/crawljax/crawljax  (Sunman et al., 2022)
Cucumber Open-source It is a tool that automatically runs business case-driven tests, ensuring that all scenarios are tested as per the acceptance criteria. https://github.com/cucumber/cucumber  (Rathod et al., 2015)
Nikto Open-source It is an open-source command line tool that focuses on quickly scanning web server and website configuration errors, known security vulnerabilities, and missing XSS protection headers. https://cirt.net/Nikto2  (Devi and Kumar, 2020)
Puppeteer Open-source It automates web interactions to perform testing. https://github.com/puppeteer/puppeteer  (Chen et al., 2021)
Cypress Open-source It is a JavaScript-based tool for automated end-to-end testing of modern web applications. https://github.com/cypress-io/cypress  (Mollah and van den Bos, 2023)
Axiom Open-source It is a dynamic infrastructure framework for distributed computing and testing that automates distributed penetration testing across cloud environments. https://github.com/pry0cc/axiom  (Prisadi et al., 2023)
Beautiful Soup Open-source It can be used to extract data from html and xml files. https://www.crummy.com/software/BeautifulSoup
https://git.launchpad.net/beautifulsoup
 (Paul and Tommy, 2018)
Commix Open-source It is a tool that automatically detects and exploits command injection vulnerabilities in web applications, including classic command injection, blind command injection, and other types of tools. https://github.com/commixproject/commix  (Stasinopoulos et al., 2019)
Katalon Studio Commercial (Free) It is easy to install and use, has powerful built-in reporting capabilities, and supports record and playback. https://katalon.com/  (Nguyen et al., 2021)
Cytestion Open-source It is a web-based tool for automated exploration and testing of GUIs. https://gitlab.com/lsi-ufcg/cytestion/cytestion  (Moura et al., 2023)
Lazyrecon Open-source It is an automated vulnerability scanning and data collection framework. https://github.com/nahamsec/lazyrecon/blob/master/lazyrecon.sh  (Shaji and Subramanian, 2021)
Jena API Open-source It is an open-source Java framework for building semantic web and linked data applications, supporting SPARQL queries. https://jena.apache.org/  (Durai et al., 2021)
Binwalk Open-source It is a firmware analysis tool suitable for most firmware unpacking. https://github.com/ReFirmLabs/binwalk  (Li et al., 2023)
jÄk Open-source It is a web application crawler and scanner tool that uses dynamic Java Script code analysis. https://github.com/ConstantinT/jAEk  (Pellegrino et al., 2015)
Skipfish Open-source It is an efficient web application security vulnerability scanning tool that can quickly detect common web application security issues. https://code.google.com/archive/p/skipfish/  (Touseef et al., 2019)
Sqlmap Open-source It is a tool for detecting and exploiting SQL injection vulnerabilities. https://sqlmap.org/  (Prisadi et al., 2023)
Scrapy Open-source It is a tool for quickly crawling HTML-based static web page content. https://github.com/scrapy/scrapy  (Rennhard et al., 2022)
Metasploit Open-source It is an open-source framework for penetration testing and vulnerability exploitation. https://www.metasploit.com/
https://github.com/rapid7/metasploit-framework
 (Jana and Oprea, 2019)
OWASP AppSensor Open-source It is a tool that detects suspicious application layer behavior and triggers events based on predefined sensors. https://owasp.org/www-project-appsensor/  (Shahrivar et al., 2023)
HTMLCS Open-source It is a tool for checking whether HTML code complies with WCAG standards. http://squizlabs.github.io/HTML_CodeSniffer/  (Campoverde-Molina et al., 2021)
GPTFuzzer Open-source It is a tool for generating web application firewall (WAF) bypass payloads, testing and discovering common web attacks such as SQL injection, cross-site scripting (XSS), and remote command execution (RCE) vulnerabilities. https://github.com/hongliangliang/gptfuzzer  (Liang et al., 2023)
SideeX Open-source It is a tool that provides cross-browser Web UI automated testing, supports recording and playback functions, and optimizes the test process through an automatic waiting mechanism. http://sideex.org  (Lee et al., 2018)
CICFlowMeter Open-source It is a tool that provides network flow signatures for intrusion detection and malicious behavior classification. https://github.com/ahlashkari/CICFlowMeter  (Rajić et al., 2023)
OWASP AppSensor Open-source It is a tool that detects suspicious application layer behavior and triggers events based on predefined sensors. https://owasp.org/www-project-appsensor/  (Shahrivar et al., 2023)
WebPageTest Commercial (Free) It is used to analyze web page performance, providing details such as page load time, resource load time, etc. https://www.webpagetest.org/  (Ali-Shahid and Sulaiman, 2015)
WebTest Open-source It is a tool that supports visualization of output coverage, helps testers understand covered and uncovered output parts, and optimizes test cases. https://github.com/git1997/VarAnalysis  (Nguyen et al., 2019)
GTmetrix Commercial (Free) It is used to measure web page loading speed and performance and provide detailed optimization suggestions. https://gtmetrix.com/  (Ali-Shahid and Sulaiman, 2015)
SikuliX Open-source It is an image-based testing tool that can control user input methods to interact with GUI elements, suitable for both web and non-web applications. http://sikulix.com  (Macchi et al., 2021)
Hakrawler Open-source It is a web crawler tool that collects web links and identifies injection points. https://github.com/hakluke/hakrawler  (Prisadi et al., 2023)
Axe Studio Commercial (Free) It is a tool that focuses on automated detection of web accessibility issues. https://www.deque.com/axe/  (Campoverde-Molina et al., 2021)
Cornipickle Open-source It is a declarative testing tool that allows developers to define layout errors using logical expressions and compare states of different snapshots. https://github.com/liflab/cornipickle  (Beroual et al., 2020)
Open-source tools like Selenium (Dawei et al., 2016), JMeter (Cai and Hu, 2014), and ZAP (Touseef et al., 2019) provide cross-platform compatibility, support multiple programming languages, and have received strong community support. They benefit from scalability, flexibility, and collaborative development through platforms like GitHub, where contributors can provide updates and improvements.
In contrast, commercial tools such as Katalon Studio (Nguyen et al., 2021) and GTmetrix (Ali-Shahid and Sulaiman, 2015) often prioritize enhanced usability and integrated reporting features, making them more suitable for testers and organizations that value refined interfaces and out-of-the-box functionality. However, limitations in the free versions may make them less suitable for certain users.
Nearly all of the reviewed tools are easily accessible through official websites or open-source platforms. This enables a smooth integration into a wide range of testing environments. Comprehensive documentation and active community support further simplify adoption and customization. The tools offer a range of diverse capabilities: From identifying security vulnerabilities with tools like Wapiti (Anagandula and Zavarsky, 2020) and Skipfish (Touseef et al., 2019), to enhancing page performance with GTmetrix (Ali-Shahid and Sulaiman, 2015) and WebPageTest (Ali-Shahid and Sulaiman, 2015). Their wide availability and applicability underscore their role in advancing WAT practices.
9. ANSWER TO RQ6: Challenges and Future Work
This section addresses RQ6, identifying unresolved challenges in WAT, and suggesting directions for future research. We start by outlining the key challenges still affecting the field. We then discuss potential areas for future work that could help address these issues and improve WAT practices.
9.1. Challenges
Despite notable advances over recent years, a number of significant challenges to WAT remain that prevent the full realization of its potential. Here we discuss several key challenges:
  • As web applications increase in complexity, test automation faces significant challenges, particularly when scaling up to handle large test suites efficiently.
    The recent use of machine learning and containerization to manage these large-scale testing processes has helped. However, limitations persist for dynamic content and asynchronous operations (Pellegrino et al., 2015). Additionally, tool fragmentation across platforms continues to affect scalability (Pelivani and Cico, 2021).
  • Balancing test coverage and execution efficiency remain challenges in web testing. Virtual DOM coverage techniques have demonstrated the potential to enhance testing efficiency for dynamic web applications by optimizing coverage metrics. However, achieving a balance between efficiency and comprehensive coverage remains an open challenge (Zou et al., 2014). Hyper-heuristic algorithms can reduce test-case redundancy, but more research into their application is required (Balera and de Santiago Júnior, 2022).
  • Failure diagnosis can be a bottleneck for WAT. Model-based approaches can help, but false positives and false negatives remain significant issues (Alkofahi et al., 2022). Furthermore, automated repair strategies may have difficulties with the dynamic content in complex web applications, indicating a need for more robust tools.
  • Maintaining test suites for dynamic web applications can also present significant challenges: Frequent updates to UI and DOM structures often break test cases (Ryou and Ryu, 2018). Although recent work has used machine learning to adapt test scripts to changes, this remains limited when dealing with significant UI overhauls.
  • Web testing tools remain fragmented, leading to inefficiencies in end-to-end testing processes. Tools like Selenium and Katalon may excel in specific areas, but they also struggle to provide cohesive solutions for more comprehensive testing needs. Recent studies have emphasized the importance of integrating these tools for more seamless cross-platform testing (Marquardt and Buhl, 2021). However, ensuring that these tools can work together across different testing environments remains a challenge (Pelivani and Cico, 2021).
  • The lack of standardized metrics for evaluating testing effectiveness remains a significant challenge. Although various tools provide performance and security-testing capabilities, they often lack consistent benchmarks, making comparisons difficult. There is a need for unified evaluation frameworks that provide reliable metrics for test coverage and efficiency (Sönmez and Kiliç, 2021). Initiatives like WebEV are helping to standardize some aspects of test-behavior analysis, but further research is necessary  (Fuad and Sakib, 2023).
9.2. Future Work
This section outlines some potential future WAT research directions, including enhancing scalability, the integration of Large Language Models (LLMs), enhanced failure diagnosis, and the development of standardized evaluation metrics.
  • Future research should enhance the scalability of test-automation frameworks to accommodate the growing complexity of modern web applications. Machine learning models should be further optimized, reducing redundancy while maintaining coverage. The use of containerization technologies, like Docker and Kubernetes, should be expanded to better enable parallel execution and to improve resource efficiency. Integrating tools into unified frameworks will also enhance automation processes and reduce fragmentation (Pelivani and Cico, 2021).
  • LLMs represent a promising approach for automating test-case generation, particularly for web forms where contextual information is key. Initial studies (Yao et al., 2024) have shown the potential of models like GPT-4 (Sanderson, 2023) and Baichuan2 (Xiao et al., 2024). However, further research is required to address security concerns and refine the models for real-world use, including for dynamic content (Li et al., 2024a).
  • Failure-diagnosis tools are critical to the precision and reliability of WAT. Many current systems have difficulties with false positives and false negatives: Future research should therefore explore adaptive systems that can better interpret the complex interactions in web applications. Combining functional and cross-browser testing tools will also enhance overall efficiency (Alkofahi et al., 2022; Pelivani and Cico, 2021).
  • Standardized metrics for WAT-tool effectiveness are also needed. The absence of universal evaluation frameworks for coverage, efficiency, and performance limits meaningful comparisons: Consistent metrics will enable more accurate assessments and foster innovation in WAT methodologies (Sönmez and Kiliç, 2021; Fuad and Sakib, 2023).
  • Multi-agent reinforcement learning (MARL) has also shown potential for enhancing test coverage and efficiency in complex web applications  (Fan et al., 2024). Initial studies have demonstrated MARL’s effectiveness for single-agent approaches. However, further research is needed to refine inter-agent communication and data-sharing protocols, reducing redundancy and optimizing exploration across web states.
  • Large vision-language models (LVLMs), such as VETL (Wang et al., 2024), can support context-aware GUI testing by generating relevant inputs and selecting UI elements based on visual context. Although LVLMs have shown potential in managing dynamic, visually complex interfaces, future work should explore hybrid approaches that combine LVLMs with heuristic methods to improve cost-efficiency and adaptability, especially in applications with frequent content updates.
10. Conclusions
This review paper has offered a detailed analysis of advances in WAT over the past decade. The review has examined a collection of computer science papers published between January 1, 2014, and December 31, 2023. A rigorous review-research methodology was employed: Five major academic databases were systematically searched, using targeted keywords, with each retrieved paper carefully examined to ensure its relevance to the study.
Our review was guided by eight research questions, which helped organize and synthesize the data. These questions addressed essential aspects of WAT, including test-case generation and execution, evaluation metrics, and available tools. The paper provides a clear understanding of the key trends and challenges in WAT over the past decade.
Various test-case generation methods were reviewed, including those designed to handle the dynamic and asynchronous nature of modern web applications. Regarding test execution, the importance of testing across different environments to reflect real-world conditions was emphasized. Besides, metrics used to evaluate WAT effectiveness, such as test coverage and error-detection rates, were also explored, highlighting the testing process’s efficiency and quality. Additionally, available WAT tools were categorized based on their functions and accessibility, providing a practical overview for researchers and practitioners.
This paper represents a useful resource for researchers and practitioners aiming to enhance the quality, reliability, and security of web applications.
Appendices
Appendix A A comprehensive overview of articles
This section presents the finalized list of 314 research papers surveyed in this work. Each paper includes the paper title, authors, publication venue, and year of publication. The complete information is provided in the following Table LABEL:TAB:wat_articles for ease of reference.
Table A.I. Overview of Included Studies.
No. Title Authors Venue Year
1 A new algorithm for detecting SQL injection attack in web application Lounis et al. Proceedings of the 2014 Science and Information Conference (SAI’14) 2014
2 An approach to detect drive-by download by observing the web page transition behaviors Matsunaka et al. Proceedings of the 9th Asia Joint Conference on Information Security (AsiaJCIS’14) 2014
3 An invariant-based approach for detecting attacks against data in web applications Ludinard et al. Application Development and Design: Concepts, Methodologies, Tools, and Applications 2014
4 Analysis for cloud testing of web application Cai et al. Proceedings of the 2nd International Conference on Systems and Informatics (ICSAI’14) 2014
5 Automated black-box detection of access control vulnerabilities in web applications Li et al. Proceedings of the 4th ACM Conference on Data and Application Security and Privacy (SP’14) 2014
6 Data flow testing of CGI based web applications Sahu et al. Proceedings of the 2014 International Conference on Information Technology (ICIT’14) 2014
7 Innovation and evolution in integrated web application testing with TTCN-3 Stepien et al. International Journal on Software Tools for Technology Transfer 2014
8 Keyword driven automated testing framework for web application Arya et al. Proceedings of the 9th International Conference on Industrial and Information Systems (ICIIS’14) 2014
9 Leveraging existing tests in automated test generation for web applications Milani Fard et al. Proceedings of the 29th ACM/IEEE International Conference on Automated Software Engineering (ASE’14) 2014
10 Link: Exploiting the web of data to generate test inputs Mariani et al. Proceedings of the 23rd International Symposium on Software Testing and Analysis (ISSTA’14) 2014
11 LTF: A model-based load testing framework for web applications Zhou et al. Proceedings of the 14th International Conference on Quality Software (QSIC’14) 2014
12 Modeling web application for cross-browser compatibility testing Li et al. Proceedings of the 15th IEEE/ACIS International Conference on Software Engineering, Artificial Intelligence, Networking and Parallel/Distributed Computing (SNPD’14) 2014
13 Performance testing web applications on the cloud Mukherjee et al. Proceedings of the IEEE 7th International Conference on Software Testing, Verification and Validation Workshops (ICSTW’14) 2014
14 Research on web application consistency testing based on model simulation Li et al. Proceedings of the 9th International Conference on Computer Science and Education (ICCSE’14) 2014
15 Sniper: A framework for detecting attack attempts to open source web applications at the run time Almurayh Proceedings of the 15th IEEE International Conference on Information Reuse and Integration (IRI’14) 2014
16 Test as a Service: A framework for web security TaaS service in cloud environment Tung et al. Proceedings of the 2014 IEEE International Symposium on Service Oriented System Engineering (SOSE’14) 2014
17 Testing web applications by unifying fuzzy and all-Pairs techniques El Youmi et al. Proceedings of the 2014 International Conference on Multimedia Computing and Systems (ICMCS’14) 2014
18 URMG: Enhanced CBMG-based method for automatically testing web applications in the cloud Xu et al. Tsinghua Science and Technology 2014
19 A keyword-driven tool for testing web applications (KeyDriver) Gupta et al. IEEE Potentials 2014
20 An automated black box approach for web vulnerability identification and attack scenario generation Akrout et al. Journal of the Brazilian Computer Society 2014
21 Detection model for SQL injection attack: An approach for preventing a web application from the SQL injection attack Buja et al. Proceedings of the 2014 IEEE Symposium on Computer Applications and Industrial Electronics (ISCAIE’14) 2014
22 DOM-based test adequacy criteria for web applications Mirzaaghaei et al. Proceedings of the 23rd International Symposium on Software Testing and Analysis (ISSTA’14) 2014
23 Efficient execution of web navigation sequences Losada et al. World Wide Web 2014
24 Model-based test case generation for web applications Nabuco et al. Proceedings of the 14th International Conference on Computational Science and Its Applications (ICCSA’14) 2014
25 Municipal web sites accessibility and usability for blind users: Preliminary results from a pilot study Pribeanu et al. Universal Access in the Information Society 2014
26 On the applicability of combinatorial testing to web application security testing: A case study Garn et al. Proceedings of the 2014 Workshop on Joining AcadeMiA and Industry Contributions to Test Automation and Model-Based Testing (JAMAICA’14) 2014
27 Performance evaluation of web based automation testing tools Angmo et al. Proceedings of the 5th International Conference-Confluence The Next Generation Information Technology Summit (Confluence’14) 2014
28 Search-based security testing of web applications Thomé et al. Proceedings of the 7th International Workshop on Search-Based Software Testing (SBST’14) 2014
29 SSOScan: Automated testing of web applications for single sign-on vulnerabilities Zhou et al. Proceedings of the 23rd USENIX Security Symposium (SEC’14) 2014
30 SymJS: Automatic symbolic testing of JavaScript web applications Li et al. Proceedings of the 22nd ACM SIGSOFT International Symposium on Foundations of Software Engineering (FSE’14) 2014
31 Usability testing research framework: Case of handicraft web-based system Isa et al. Proceedings of the 2014 2nd International Conference on Information and Communication Technology (ICoICT’14) 2014
32 Virtual DOM coverage for effective testing of dynamic web applications Zou et al. Proceedings of the 2014 International Symposium on Software Testing and Analysis (ISSTA’14) 2014
33 A framework for testing web applications using action word based testing Garg et al. Proceedings of the 1st International Conference on Next Generation Computing Technologies (NGCT’15) 2015
34 A new framework of security vulnerabilities detection in PHP web application Zhao et al. Proceedings of the 9th International Conference on Innovative Mobile and Internet Services in Ubiquitous Computing (IMIS’15) 2015
35 A sampling method based on URL clustering for fast web accessibility evaluation Zhang et al. Frontiers of Information Technology and Electronic Engineering 2015
36 Analysis of effectiveness of black-box web application scanners in detection of stored SQL injection and stored XSS vulnerabilities Anagandula et al. Proceedings of the 3rd International Conference on Data Intelligence and Security (ICDIS’20) 2015
37 Automated testing of web applications with text input Wu et al. Proceedings of the 2015 IEEE International Conference on Progress in Informatics and Computing (PIC’15) 2015
38 Automatic performance analysis of cloud based load testing of web-application & its comparison with traditional load testing Arslan et al. Proceedings of the 6th IEEE International Conference on Software Engineering and Service Science (ICSESS’15) 2015
39 Beyond page objects: Testing web applications with state objects: Use states to drive your tests Van Deursen Queue 2015
40 Client-oriented web alteration detection system using link change state of a web page based on past and current page content Mochizuki et al. Proceedings of the 17th International Conference on Information Integration and Web-based Applications and Services (iiWAS’15) 2015
41 Evaluation of the IPO-Family algorithms for test case generation in web security testing Bozic et al. Proceedings of the IEEE 8th International Conference on Software Testing, Verification and Validation Workshops (ICSTW’15) 2015
42 Event-driven web application testing based on model-based mutation testing Habibi et al. Information and Software Technology 2015
43 Meta-heuristic generation of robust XPath locators for web testing Leotta et al. Proceedings of the IEEE/ACM 8th International Workshop on Search-Based Software Testing (SBST’15) 2015
44 Performance testing as a service for web applications Ali et al. Proceedings of the IEEE 7th International Conference on Intelligent Computing and Information Systems (ICICIS’15) 2015
45 SQLi penetration testing of financial web applications: Investigation of Bangladesh region Farah et al. Proceedings of the 2015 World Congress on Internet Security (WorldCIS’15) 2015
46 Testing web applications with state objects Van Deursen Communications of the ACM 2015
47 Text-mining based predictive model to detect XSS vulnerable files in web applications Gupta et al. Proceedings of the 2015 Annual IEEE India Conference (INDICON’15) 2015
48 Usability testing on government agencies web portal: A study on Ministry of Education Malaysia (MOE) web portal Rozali et al. Proceedings of the 2015 9th Malaysian Software Engineering Conference (MySEC’15) 2015
49 Vulnerability assessment of web applications - A testing approach Vibhandik et al. Proceedings of the 4th International Conference on e-Technologies and Networks for Development (ICeND’15) 2015
50 Web from preprocessor for crawling Muñoz et al. Multimedia Tools and Applications 2015
51 A case study on reliability and usability testing of a web portal Ali-Shahid et al. Proceedings of the 9th Malaysian Software Engineering Conference (MySEC’15) 2015
52 A novel anomaly detection approach for mitigating web-based attacks against clouds Zhang et al. Proceedings of the IEEE 2nd International Conference on Cyber Security and Cloud Computing (CSCloud’15) 2015
53 An extensible environment for guideline-based accessibility evaluation of dynamic web applications Schiavone et al. Universal Access in the Information Society 2015
54 Attack pattern-based combinatorial testing with constraints for web security testing Bozic et al. Proceedings of the 2015 IEEE International Conference on Software Quality, Reliability and Security (QRS’15) 2015
55 Automatic acceptance testing of the web application security with ITU-T X.805 framework Rathod et al. Proceedings of the 2nd International Conference on Computer Science, Computer Engineering, and Social Media (CSCESM’15) 2015
56 Automatic generation of valid and invalid test data for string validation routines using web searches and regular expressions Shahbaz et al. Science of Computer Programming 2015
57 Detecting and exploiting second order denial-of-service vulnerabilities in web applications Olivo et al. Proceedings of the 22nd ACM SIGSAC Conference on Computer and Communications Security (CCS’15) 2015
58 Detection of web appication vulnerability based on RUP model Gol et al. Proceedings of the 2015 National Conference on Recent Advances in Electronics & Computer Engineering (RAECE’15) 2015
59 ER diagram based web application testing Begum et al. Proceedings of the 2015 IEEE International Conference on Computational Intelligence and Computing Research (ICCIC’15) 2015
60 Experiences in performance testing of web applications with unified authentication platform using Jmeter Kiran et al. Proceedings of the 2015 International Symposium on Technology Management and Emerging Technologies (ISTMET’15) 2015
61 Guided mutation testing for JavaScript web applications Mirshokraie et al. IEEE Transactions on Software Engineering 2015
62 Heuristic evaluation as a complement to usability testing: A case study in web domain Paz et al. Proceedings of the 2015 12th international conference on information technology-new generations (ITNG’15) 2015
63 Incremental web application testing using page object Yu et al. Proceedings of the 3rd IEEE Workshop on Hot Topics in Web Systems and Technologies (HotWeb’15) 2015
64 jÄk: Using dynamic analysis to crawl and test modern web applications Pellegrino et al. Proceedings of the 18th International Symposium on Research in Attacks, Intrusions, and Defenses (RAID’15) 2015
65 Leveraging task-based data to support functional testing of web applications de Jesus et al. Proceedings of the 30th Annual ACM Symposium on Applied Computing (SAC’15) 2015
66 Orchestration framework for automated Ajax-based web application testing Deyab et al. Proceedings of the 9th Malaysian Software Engineering Conference (MySEC’15) 2015
67 Syntax-based test case generation for web application Polpong et al. Proceedings of the 2015 International Conference on Computer, Communications, and Control Technology (I4CT’15) 2015
68 A grey-box approach for detecting malicious user interactions in web applications Ben Jaballah et al. Proceedings of the 8th ACM CCS International Workshop on Managing Insider Security Threats (MIST’16) 2016
69 A signature-based intrusion detection system for web applications based on genetic algorithm Bronte et al. Proceedings of the 9th International Conference on Security of Information and Networks (SIN’16) 2016
70 An automated approach for testing the security of web applications against chained attacks Calvi et al. Proceedings of the 31st Annual ACM Symposium on Applied Computing (SAC’16) 2016
71 An executable model and testing for web software based on live sequence charts Li et al. Proceedings of the 2016 IEEE/ACIS 15th International Conference on Computer and Information Science (ICIS’16) 2016
72 Automated white-list learning technique for detection of malicious attack on web application Murtaza et al. Proceedings of the 13th International Bhurban Conference on Applied Sciences and Technology (IBCAST’16) 2016
73 Composing and delivering heterogeneous web testing software as a composite web testing service Lee et al. Proceedings of the 2016 International Computer Symposium (ICS’16) 2016
74 Data leakage detection in Tizen web applications Rocha et al. Proceeding of the 14th Annual Conference on Privacy, Security and Trust (PST’16) 2016
75 DEKANT: A static analysis tool that learns to detect web application vulnerabilities Medeiros et al. Proceedings of the 25th International Symposium on Software Testing and Analysis (ISSTA’16) 2016
76 Detecting privilege escalation attacks through instrumenting web application source code Zhu et al. Proceedings of the 21st ACM Symposium on Access Control Models and Technologies (SACMAT’16) 2016
77 Detecting web attacks using multi-stage log analysis Moh et al. Proceedings of the IEEE 6th international conference on advanced computing (IACC’16) 2016
78 Web application automatic testing solution Dawei et al. Proceedings of the 3rd International Conference on Information Science and Control Engineering (ICISCE’16) 2016
79 Web application’s performance testing using HP LoadRunner and CA Wily introscope tools Khan et al. Proceedings of the 2016 International Conference on Computing, Communication and Automation (ICCCA’16) 2016
80 WePR: A tool for automated web performance measurement Asrese et al. Proceedings of the 2016 IEEE Globecom Workshops (GC Wkshps’16) 2016
81 Detect cross-browser issues for JavaScript-based web applications based on record/replay Wu et al. Proceedings of the 32nd IEEE International Conference on Software Maintenance and Evolution (ICSME’16) 2016
82 End to end specification based test generation of web applications Bubna et al. Proceedings of the 11th International Conference on Evaluation of Novel Software Approaches to Software Engineering (ENASE’16) 2016
83 Fragmented query parse tree based SQL injection detection system for web applications Priyaa et al. Proceedings of the 2016 International Conference on Computing Technologies and Intelligent Data Engineering (ICCTIDE’16) 2016
84 Implementation of extension for browser to detect vulnerable elements on web pages and avoid Clickjacking Pawade et al. Proceedings of the 2016 6th International Conference-Cloud System and Big Data Engineering (Confluence’16) 2016
85 Smoke testing of web application based on ALM tool Khan et al. Proceedings of the 2016 International Conference on Computing, Communication and Automation (ICCCA’16) 2016
86 Static DOM event dependency analysis for testing web applications Sung et al. Proceedings of the 2016 24th ACM SIGSOFT International Symposium on Foundations of Software Engineering (FSE’16) 2016
87 Testing of Ajax-based web applications using hierarchical state model Liu et al. Proceedings of the 2016 IEEE 13th International Conference on e-Business Engineering (ICEBE’16) 2016
88 The method of generating web link security testing scenario based on UML diagram Wang et al. Proceedings of the 15th IEEE International Conference on Trust, Security and Privacy in Computing and Communications (TrustCom’16) 2016
89 Using combinatorial testing to build navigation graphs for dynamic web applications Wang et al. Software Testing, Verification and Reliability 2016
90 Using Petri nets to test concurrent behavior of web applications Thummala et al. Proceedings of the IEEE 9th International Conference on Software Testing, Verification and Validation Workshops (ICSTW’16) 2016
91 Web database testing using ER diagram and state transition model Indumathi et al. Proceedings of the 2016 International Conference on Communication and Signal Processing (ICCSP’16) 2016
92 WebGuardia - An integrated penetration testing system to detect web application vulnerabilities Vithanage et al. Proceedings of the 2016 International Conference on Wireless Communications, Signal Processing and Networking (WiSPNET’16) 2016
93 A study on acceptance procedure improvement of web application by outsourcing for mobile service Choi et al. Wireless Personal Communications 2017
94 An approach for dynamic web application testing using MBT Panthi et al. International Journal of System Assurance Engineering and Management 2017
95 Universities of the Kyrgyz Republic on the web: Accessibility and usability Ismailova et al. Universal Access in the Information Society 2017
96 Using semantic similarity in crawling-based web application testing Lin et al. Proceedings of the 10th IEEE International Conference on Software Testing, Verification and Validation (ICST’17) 2017
97 Web site accessibility and quality in use: A comparative study of government web sites in Kyrgyzstan, Azerbaijan, Kazakhstan and Turkey Ismailova et al. Universal Access in the Information Society 2017
98 A search-based testing approach for XML injection vulnerabilities in web applications Jan et al. Proceedings of the 10th IEEE International Conference on Software Testing, Verification and Validation (ICST’17) 2017
99 An anomaly detection method based on multi-models to detect web attacks Zhang et al. Proceedings of the 10th International Symposium on Computational Intelligence and Design (ISCID’17) 2017
100 Automated complete test case coverage for web based applications Nagowah et al. Proceedings of the 2017 International Conference on Infocom Technologies and Unmanned Systems (Trends and Future Directions) (ICTUS’17) 2017
101 Automated testing of web applications using combinatorial strategies Qi et al. Journal of Computer Science and Technology 2017
102 Customizable automatic detection of bad usability smells in mobile accessed web applications Paternò et al. Proceedings of the 19th International Conference on Human-Computer Interaction with Mobile Devices and Services (MobileHCI’17) 2017
103 Detection of malicious web pages based on hybrid analysis Wang et al. Journal of Information Security and Applications 2017
104 Evaluation of deception-based web attacks detection Han et al. Proceedings of the 2017 Workshop on Moving Target Defense (MTD’17) 2017
105 Feature driven hybrid test automation framework (FDHTAF) for web based or cloud based application testing Kiranagi et al. Proceedings of the 2017 International Conference On Smart Technologies For Smart Nation (SmartTechCon’17) 2017
106 Lom: Discovering logic flaws within MongoDB-based web applications Wen et al. International Journal of Automation and Computing 2017
107 On the usage of character distribution for the detection of web attacks Le Dinh et al. Proceedings of the 2017 9th International Conference on Knowledge and Systems Engineering (KSE’17) 2017
108 Parameter manipulation attack prevention and detection by using web application deception proxy Ishikawa et al. Proceedings of the 11th international conference on ubiquitous information management and communication (IMCOM’17) 2017
109 Performance testing analysis on web application: Study case student admission web system Putri et al. Proceedings of the 2017 International Conference on Sustainable Information Engineering and Technology (SIET’17) 2017
110 Searching for behavioural bugs with stateful test oracles in web crawlers Beroual et al. Proceedings of the IEEE/ACM 10th International Workshop on Search-Based Software Testing (SBST’17) 2017
111 Test case generation from BPMN models for automated testing of web-based BPM applications de Moura et al. Proceedings of the 17th International Conference on Computational Science and Its Applications (ICCSA’17) 2017
112 Test execution checkpointing for web applications Guarnieri et al. Proceedings of the 26th ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA’17) 2017
113 Testing and analysis of web applications using page models Athaiya et al. Proceedings of the 26th ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA’17) 2017
114 Towards continuous security certification of software-as-a-service applications using web application testing techniques Stephanow et al. Proceedings of the 2017 IEEE 31st International Conference on Advanced Information Networking and Applications (AINA’17) 2017
115 Towards the generation of end-to-end web test scripts from requirements specifications Clerissi et al. Proceedings of the 2017 IEEE 25th International Requirements Engineering Conference Workshops (REW’17) 2017
116 User acceptance testing for Agile-developed web-based applications: Empowering customers through wikis and mind maps Otaduy et al. Journal of Systems and Software 2017
117 Vulnerability assessment and penetration testing of web application Nagpure et al. Proceedings of the 2017 International Conference on Computing, Communication, Control and Automation (ICCUBEA’17) 2017
118 Web application fuzz testing Andrianto et al. Proceedings of the 2017 International Conference on Data and Software Engineering (ICoDSE’17) 2017
119 Web site accessibility, usability and security: A survey of government web sites in Kyrgyz Republic Ismailova Universal Access in the Information Society 2017
120 An algorithm to find relationships between web vulnerabilities Román Muñoz et al. The Journal of Supercomputing 2018
121 Black-box detection of XQuery injection and parameter tampering vulnerabilities in web applications Deepa et al. International Journal of Information Security 2018
122 Checking compliance of semantic web applications with RDFS-semantics Colucci et al. Internet Technology Letters 2018
123 HoneyCirculator: Distributing credential honeytoken for introspection of web-based attack cycle Akiyama et al. International Journal of Information Security 2018
124 A machine-learning-driven evolutionary approach for testing web application firewalls Appelt et al. IEEE Transactions on Reliability 2018
125 A web second-order vulnerabilities detection method Liu et al. IEEE Access 2018
126 An approach of automated testing on web based platform using machine learning and Selenium Paul et al. Proceedings of the 2018 International Conference on Inventive Research in Computing Applications (ICIRCA’18) 2018
127 Analytical study and implementation of web performance testing tools Saharan et al. Proceedings of the 2018 International Conference on Recent Innovations in Electrical, Electronics and Communication Engineering (ICRIEECE’18) 2018
128 Automatic detection of visibility faults by layout changes in HTML5 web pages Ryou et al. Proceedings of the IEEE 11th International Conference on Software Testing, Verification and Validation (ICST’18) 2018
129 Automating web application testing from the ground up: Experiences and lessons learned in an industrial setting Debroy et al. Proceedings of the IEEE 11th International Conference on Software Testing, Verification and Validation (ICST’18) 2018
130 Capability matching and heuristic search for job assignment in crowdsourced web application testing Guo et al. Proceedings of the 2018 IEEE International Conference on Systems, Man, and Cybernetics (SMC’18) 2018
131 Cross-browser differences detection based on an empirical metric for web page visual similarity Xu et al. ACM Transactions on Internet Technology 2018
132 Detecting attacks on web applications using autoencoder Mac et al. Proceedings of the 9th International Symposium on Information and Communication Technology (SoICT’18) 2018
133 DetLogic: A black-box approach for detecting logic vulnerabilities in web applications Deepa et al. Journal of Network and Computer Applications 2018
134 Enlargement of vulnerable web applications for testing Román Muñoz et al. The Journal of Supercomputing 2018
135 Fuzzing process modelling for web applications robustness testing via dynamic Bayesian networks Azarnova et al. Proceedings of the 2018 Global Smart Industry Conference (GloSIC’18) 2018
136 Identifying worst-case user scenarios for performance testing of web applications using Markov-chain workload models Ahmad et al. Future Generation Computer Systems 2018
137 Malicious hidden redirect attack web page detection based on CSS features Chen et al. Proceedings of the 2018 IEEE 4th International Conference on Computer and Communications (ICCC’18) 2018
138 MobSTer: A model-based security testing framework for web applications Peroli et al. Software Testing, Verification and Reliability 2018
139 Pattern-based web security testing Araújo et al. Proceedings of the 6th International Conference on Model-Driven Engineering and Software Development (MODELSWARD’18) 2018
140 Planning-based security testing of web applications Bozic et al. Proceedings of the 13th International Workshop on Automation of Software Test (AST’18) 2018
141 Studying the effectiveness of various tools in detecting the protecting mechanisms implemented in web-applications Thakre et al. Proceedings of the 2018 International Conference on Inventive Research in Computing Applications (ICIRCA’18) 2018
142 Test case generation for semantic-based user input validation of web applications Hanna et al. International Journal of Web Engineering and Technology 2018
143 Test command auto-wait mechanisms for record and playback-style web application testing Lee et al. Proceedings of the IEEE 42nd Annual Computer Software and Applications Conference (COMPSAC’18) 2018
144 Tools for active and passive network side-channel detection for web applications Lescisin et al. Proceedings of the 12th USENIX Workshop on Offensive Technologies (WOOT’18) 2018
145 Using the visual method of HTTPS checking for measuring the level of data transport security on web-based smart cities applications in Indonesia Wibowo Proceedings of the 2018 International Conference on ICT for Smart Society (ICISS’18) 2018
146 Vesper: Measuring time-to-interactivity for web pages Netravali et al. Proceedings of the 15th USENIX Symposium on Networked Systems Design and Implementation (NSDI’18) 2018
147 Vulnerability coverage criteria for security testing of web applications Murthy et al. Proceedings of the 2018 international conference on advances in computing, communications and informatics (ICACCI’18) 2018
148 Vulnerability detection in PHP web application using lexical analysis approach with machine learning Anbiya et al. Proceedings of the 2018 5th International Conference on Data and Software Engineering (ICoDSE’18) 2018
149 An empirical study on combining diverse static analysis tools for web security vulnerabilities based on development scenarios Nunes et al. Computing 2019
150 Automated combinatorial testing for detecting SQL vulnerabilities in web applications Simos et al. Proceedings of the IEEE/ACM 14th International Workshop on Automation of Software Test (AST’19) 2019
151 Commix: Automating evaluation and exploitation of command injection vulnerabilities in Web applications Stasinopoulos et al. International Journal of Information Security 2019
152 Design and implementation of core modules of web application vulnerability detection model Ze Proceedings of the 11th International Conference on Measuring Technology and Mechatronics Automation (ICMTMA’19) 2019
153 Knowledge-based security testing of web applications by logic programming Zech et al. International Journal on Software Tools for Technology Transfer 2019
154 Test case generation based on client-server of web applications by memetic algorithm Wang et al. Proceedings of the IEEE 30th International Symposium on Software Reliability Engineering (ISSRE’19) 2019
155 Testing concurrent user behavior of synchronous web applications with Petri nets Offutt et al. Software and Systems Modeling 2019
156 A pragmatic evaluation of stress and performance testing technologies for web based applications Pradeep et al. Proceedings of the 2019 Amity International Conference on Artificial Intelligence (AICAI’19) 2019
157 An approach for web applications test data generation based on analyzing client side user input fields Hanna et al. Proceedings of the 2nd International Conference on New Trends in Computing Sciences (ICTCS’19) 2019
158 Analysis of automated web application security vulnerabilities testing Touseef et al. Proceedings of the 3rd International Conference on Future Networks and Distributed Systems (ICFNDS’19) 2019
159 AppMine: Behavioral analytics for web application vulnerability detection Jana et al. Proceedings of the 2019 ACM SIGSAC Conference on Cloud Computing Security Workshop (CCSW’19) 2019
160 Automated model-based test case generation for web user interfaces (WUI) from interaction flow modeling language (IFML) models Yousaf et al. IEEE Access 2019
161 Convolutional neural network with character embeddings for malicious web request detection Wu et al. Proceedings of the 2019 IEEE Intl Conf on Parallel & Distributed Processing with Applications, Big Data & Cloud Computing, Sustainable Computing & Communications, Social Computing & Networking (ISPA/BDCloud/SocialCom/SustainCom’19) 2019
162 DeepWAF: Detecting web attacks based on CNN and LSTM models Kuang et al. Proceedings of the 11th International Symposium Cyberspace Safety and Security (CSS’19) 2019
163 Exploring output-based coverage for testing PHP web applications Nguyen et al. Automated Software Engineering 2019
164 Generating mock skeletons for lightweight web-service testing Bhagya et al. Proceedings of the 26th Asia-Pacific Software Engineering Conference (APSEC’19) 2019
165 Implementation of blockchain based distributed web attack detection application Tanrıverdi et al. Proceedings of the 1st International Informatics and Software Engineering Conference (UBMYK’19) 2019
166 JCOMIX: A search-based tool to detect XML injection vulnerabilities in web applications Stallenberg et al. Proceedings of the 2019 27th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE’19) 2019
167 Kaleidoscope: A crowdsourcing testing tool for web quality of experience Wang et al. Proceedings of the 2019 IEEE 39th International Conference on Distributed Computing Systems (ICDCS’19) 2019
168 Leveraging keyword-guided exploration to build test models for web applications Qi et al. Information and Software Technology 2019
169 Measuring web latency and rendering performance: Method, tools, and longitudinal dataset Asrese et al. IEEE Transactions on Network and Service Management 2019
170 Method for generating test data for detecting SQL injection vulnerability in web application Awang et al. Proceedings of the 7th International Conference on Cyber and IT Service Management (CITSM’19) 2019
171 MLPXSS: An integrated XSS-based attack detection scheme in web applications using multilayer perceptron technique Mokbal et al. IEEE Access 2019
172 PathMarker: Protecting web contents against inside crawlers Wan et al. Cybersecurity 2019
173 Scenario testing of AngularJS-based single page web applications Zhang et al. International Conference on Web Engineering 2019
174 Search-based multi-vulnerability testing of XML injections in web applications Jan et al. Empirical Software Engineering 2019
175 SQL injection detection for web applications based on elastic-pooling CNN Xie et al. IEEE Access 2019
176 Static code analysis tools with the taint analysis method for detecting web application vulnerability Maskur et al. Proceedings of the 2019 International Conference on Data and Software Engineering (ICoDSE’19) 2019
177 Test them all, is it worth it? Assessing configuration sampling on the JHipster web development stack Halin et al. Empirical Software Engineering 2019
178 Testing for integrity flaws in web sessions Calzavara et al. Proceedings of the Computer Security–ESORICS 2019: 24th European Symposium on Research in Computer Security (ESORICS’19) 2019
179 Web test dependency detection Biagiola et al. Proceedings of the 2019 27th ACM Joint Meeting on European Software Engineering Conference and Symposium on the Foundations of Software Engineering (ESEC/FSE’19) 2019
180 Web view: Measuring & monitoring representative information on websites Saverimoutou et al. Proceedings of the 2019 22nd Conference on Innovation in Clouds, Internet and Networks and Workshops (ICIN’19) 2019
181 WebHound: A data-driven intrusion detection from real-world web access logs Wei et al. Soft Computing 2019
182 What methods software teams prefer when testing web accessibility Bai et al. Advances in Human-Computer Interaction 2019
183 An automated framework for evaluating open-source web scanner vulnerability severity Amankwah et al. Service Oriented Computing and Applications 2020
184 Detecting responsive web design bugs with declarative specifications Beroual et al. Proceedings of the 20th International Conference on Web Engineering (ICWE’20) 2020
185 Detection of structure query language injection vulnerability in web-driven database application Aliero et al. Concurrency and Computation: Practice and Experience 2020
186 GUIDE: An interactive and incremental approach for crawling web applications Liu et al. The Journal of Supercomputing 2020
187 Metamorphic security testing for web systems Mai et al. Proceedings of the IEEE 13th International Conference on Software Testing, Validation and Verification (ICST’20) 2020
188 Planning-based security testing of web applications with attack grammars Bozic et al. Software quality journal 2020
189 Zerowall: Detecting zero-day web attacks through encoder-decoder recurrent neural networks Tang et al. Proceedings of the IEEE INFOCOM 2020 – IEEE Conference on Computer Communications (INFOCOM’20) 2020
190 A family of experiments to assess the impact of page object pattern in web test suite development Leotta et al. Proceedings of the IEEE 13th International Conference on Software Testing, Validation and Verification (ICST’20) 2020
191 An analysis of effectiveness of black-box web application scanners in detection of stored SQL injection and stored XSS vulnerabilities Anagandula et al. Proceedings of the 3rd International Conference on Data Intelligence and Security (ICDIS’20) 2020
192 Automated versus manual approach of web application penetration testing Singh et al. Proceedings of the 11th International Conference on Computing, Communication and Networking Technologies (ICCCNT’20) 2020
193 Automatic detection of usability smells in web applications running in mobile devices Limaylla Lunarejo et al. Proceedings of the XVI Brazilian Symposium on Information Systems (SBSI’20) 2020
194 Detection of XSS in web applications using machine learning classifiers Banerjee et al. Proceedings of the 4th International Conference on Electronics, Materials Engineering and Nano-Technology (IEMENTech’20) 2020
195 Effect of coding styles in detection of web application vulnerabilities Medeiros et al. Proceedings of the 16th European Dependable Computing Conference (EDCC’20) 2020
196 Fostering the diversity of exploratory testing in web applications Leveau et al. Software Testing, Verification and Reliability 2020
197 Generation of realistic navigation paths for web site testing using recurrent neural networks and generative adversarial neural networks Pavanetto et al. Proceedings of the International Conference on Web Engineering (ICWE’20) 2020
198 Machine learning for web vulnerability detection: The case of cross-site request forgery Calzavara et al. IEEE Security and Privacy 2020
199 Maintainability of automatic acceptance tests for web applications — A case study comparing two approaches to organizing code of test cases Sadaj et al. Proceedings of the 46th International Conference on Current Trends in Theory and Practice of Informatics (SOFSEM’20) 2020
200 MERLIN: Multi-language web vulnerability detection Figueiredo et al. Proceedings of the IEEE 19th International Symposium on Network Computing and Applications (NCA’20) 2020
201 Model driven approach for automatic script generation in stress testing of web applications Waheed et al. Proceedings of the 6th International Conference on Computer and Technology Applications (ICCTA’20) 2020
202 Near-duplicate detection in web app model inference Yandrapally et al. Proceedings of the ACM/IEEE 42nd International Conference on Software Engineering (ICSE’20) 2020
203 New biostatistics features for detecting web bot activity on web applications Rahman et al. Computers and Security 2020
204 Ontology-driven security testing of web applications Bozic et al. Proceedings of the 2020 IEEE International Conference on Artificial Intelligence Testing (AITest’20) 2020
205 Performance-based comparative analysis of open source vulnerability testing tools for web database applications Mishra et al. Proceedings of the 11th International Conference on Computing, Communication and Networking Technologies (ICCCNT’20) 2020
206 Research on scoring method of web application vulnerability scanning product performance test Wu et al. Proceedings of the 2020 4th Annual International Conference on Data Science and Business Analytics (ICDSBA’20) 2020
207 Testing for security weakness of web applications using ethical hacking Devi et al. Proceedings of the 2020 4th International Conference on Trends in Electronics and Informatics (ICOEI’20) 2020
208 Towards a deep learning model for vulnerability detection on web application variants Fidalgo et al. Proceedings of the 13th IEEE International Conference on Software Testing, Verification and Validation Workshops (ICSTW’20) 2020
209 Web accessibility of healthcare web sites of Korean government and public agencies: A user test for persons with visual impairment Yi Universal Access in the Information Society 2020
210 Web accessibility testing for deaf: Requirements and approaches for automation Sousa et al. Proceedings of the 2020 IEEE International Conference on Systems, Man, and Cybernetics (SMC’20) 2020
211 Web accessibility testing for Singapore government e-services Lim et al. Proceedings of the 17th International Web for All Conference (W4A’20) 2020
212 Web application testing with model based testing method: Case study Akpinar et al. Proceedings of the 2020 International Conference on Electrical, Communication, and Computer Engineering (ICECCE’20) 2020
213 Wsad: An unsupervised web session anomaly detection method Sun et al. Proceedings of the 2020 16th International Conference on Mobility, Sensing and Networking (MSN’20) 2020
214 Analyzing and eliminating phishing threats in IoT, network and other web applications using iterative intersection Nirmal et al. Peer-to-Peer Networking and Applications 2021
215 Defences against web application attacks and detecting phishing links using machine learning Hashim et al. Proceedings of the 2020 International Conference on Computer, Control, Electrical, and Electronics Engineering (ICCCEEE’21) 2021
216 Generating test scripts for web-based applications Sriwichainan et al. Proceedings of the 3rd International Conference on Sensors, Signal and Image Processing (SSIP’20) 2021
217 Image-based approaches for automating GUI testing of interactive web-based applications Macchi et al. Proceedings of the 28th Conference of Open Innovations Association (FRUCT’21) 2021
218 Model-based automated testing of JavaScript web applications via longer test sequences Gao et al. Frontiers of Computer Science 2021
219 A comparative study of automation testing tools for web applications Pelivani et al. Proceedings of the 10th Mediterranean Conference on Embedded Computing (MECO’21) 2021
220 A comparative study of RNN-based methods for web malicious code detection Guan et al. Proceedings of the IEEE 6th International Conference on Computer and Communication Systems (ICCCS’21) 2021
221 A novel method to detect and prevent SQLIA using ontology to cloud web security Durai et al. Wireless Personal Communications 2021
222 Assessing non-intrusive vulnerability scanning methodologies for detecting web application vulnerabilities on large scale Shaji et al. Proceedings of the 2021 International Conference on System, Computation, Automation and Networking (ICSCAN’21) 2021
223 Automated test data generation for Typescript web applications Thu et al. Proceedings of the 13th International Conference on Knowledge and Systems Engineering (KSE’21) 2021
224 Automatic detection and risk assessment of session management vulnerabilities in web applications Garmabi et al. Proceedings of the 11th International Conference on Computer Engineering and Knowledge (ICCKE’21) 2021
225 Automatic web testing using curiosity-driven reinforcement learning Zheng et al. Proceedings of the IEEE/ACM 43rd International Conference on Software Engineering (ICSE’21) 2021
226 Automatic web-based software structural testing using an adaptive particle swarm optimization algorithm for test data generation Damia et al. Proceedings of the 7th International Conference on Web Research (ICWR’21) 2021
227 Déjà Vu? Client-side fingerprinting and version detection of web application software Marquardt et al. Proceedings of the IEEE 46th Conference on Local Computer Networks (LCN’21) 2021
228 Detecting indicators of compromise in web applications using access logs Cernica et al. Proceedings of the 2021 IEEE International Black Sea Conference on Communications and Networking (BlackSeaCom’21) 2021
229 Detection of advanced web bots by combining web logs with mouse behavioural biometrics Iliou et al. Digital Threats: Research and Practice 2021
230 Development of a web-resources testing system for compliance with GDPR regulation Amanzholova et al. Proceedings of the 7th International Conference on Engineering & MIS (ICEMIS’21) 2021
231 Elastic performance test method of web server in cloud computing environment Su et al. Journal of Web Engineering 2021
232 Generating and selecting resilient and maintainable locators for web automated testing Nguyen et al. Software Testing, Verification and Reliability 2021
233 Generation of realistic navigation paths for web site testing using RNN and GAN Pavanetto et al. Journal of Web Engineering 2021
234 Guide of principles and good practices for software security testing in web applications for a private sector company Bautista et al. Proceedings of the 2021 Congreso Internacional de Innovación y Tendencias en Ingeniería (CONIITI’21) 2021
235 Holistic web application security visualization for multi-project and multi-phase dynamic application security test results Sönmez et al. IEEE Access 2021
236 JSON web token penetration testing on cookie storage with CSRF techniques Darmawan et al. Proceedings of the 2021 International Conference Advancement in Data Science, E-learning and Information Systems (ICADEIS’21) 2021
237 Measuring web session security at scale Calzavara et al. Computers and Security 2021
238 Model-based testing in practice: An experience report from the web applications domain Garousi et al. Journal of Systems and Software 2021
239 NLP-assisted web element identification toward script-free testing Kirinuki et al. Proceedings of the 37th IEEE International Conference on Software Maintenance and Evolution (ICSME’21) 2021
240 Novel detection technique for framejacking vulnerabilities in web applications Kalim et al. Proceedings of the 2021 2nd International Conference on Computation, Automation and Knowledge Management (ICCAKM’21) 2021
241 Process model for continuous testing of web accessibility Campoverde-Molina et al. IEEE Access 2021
242 Risk assessment of web application penetration testing on cross-site request forgery (CSRF) attacks and server-side includes (SSI) injections Zakaria et al. Proceedings of the 2021 International Conference on Data Science and Its Applications (ICoDSA’21) 2021
243 Semantic web accessibility testing via hierarchical visual analysis Bajammal et al. Proceedings of the 2021 IEEE/ACM 43rd International Conference on Software Engineering (ICSE’21) 2021
244 Sidereal: Statistical adaptive generation of robust locators for web testing Leotta et al. Software Testing, Verification and Reliability 2021
245 STILE: A tool for parallel execution of E2E web test scripts Olianas et al. Proceedings of the 14th IEEE Conference on Software Testing, Verification and Validation (ICST’21) 2021
246 Test driven development in action: case study of a cross-platform web application Pombo et al. Proceedings of the IEEE EUROCON 2021-19th International Conference on Smart Technologies (EUROCON’21) 2021
247 Using deep learning to construct auto web penetration test Jiao et al. Proceedings of the 13th International Conference on Machine Learning and Computing (ICMLC’21) 2021
248 Using dynamic analysis to automatically detect anti-adblocker on the web Sun et al. Proceedings of the IEEE INFOCOM 2021-IEEE Conference on Computer Communications Workshops (INFOCOM WKSHPS’21) 2021
249 Web application attacks detection using deep learning Montes et al. Proceedings of the 25th Iberoamerican Congress on Progress in Pattern Recognition, Image Analysis, Computer Vision, and Applications (CIARP’21) 2021
250 Web application testing: Using tree kernels to detect near-duplicate states in automated model inference Corazza et al. Proceedings of the 15th ACM/IEEE International Symposium on Empirical Software Engineering and Measurement (ESEM’21) 2021
251 X-Check: Improving effectiveness and efficiency of cross-browser issues detection for JavaScript-based web applications Wu et al. IEEE Transactions on Services Computing 2021
252 Cyber attack detection and implementation of prevention methods for web application Bhalme et al. Proceedings of the 2022 IEEE Bombay Section Signature Conference (IBSSC’22) 2022
253 Model-based testing leveraged for automated web tests Mattiello et al. Software Quality Journal 2022
254 RAT: Reinforcement-learning-driven and adaptive testing for vulnerability discovery in web application firewalls Amouei et al. IEEE Transactions on Dependable and Secure Computing 2022
255 The agent web model: Modeling web hacking for reinforcement learning Erdődi et al. International Journal of Information Security 2022
256 An enhanced deep learning based framework for web attacks detection, mitigation and attacker profiling Shahid et al. Journal of Network and Computer Applications 2022
257 Automated web application testing driven by pre-recorded test cases Sunman et al. Journal of Systems and Software 2022
258 Automating the detection of access control vulnerabilities in web applications Rennhard et al. SN Computer Science 2022
259 Beaconnect: Continuous web performance A/B testing at scale Wingerath et al. Proceedings of the VLDB Endowment 2022
260 Design and implementation of system of the web vulnerability detection based on crawler and natural language processing Ge et al. Proceedings of the 7th International Conference on Information and Network Technologies (ICINT’22) 2022
261 Design and Implementation of web application penetration testing with cross-site scripting (XSS) Goto et al. Proceedings of the IEEE 2nd International Conference on Mobile Networks and Wireless Communications (ICMNWC’22) 2022
262 Discovering authorization business rules toward detecting web applications logic flaws Alkofahi et al. Proceedings of the 2022 International Arab Conference on Information Technology (ACIT’22) 2022
263 Early web application attack detection using network traffic analysis Rajić et al. International Journal of Information Security 2022
264 eBAT: An efficient automated web application testing approach based on tester’s behavior Fuad et al. Proceedings of the 2022 29th Asia-Pacific Software Engineering Conference (APSEC’22) 2022
265 Evaluating the potential implementation of new metrics in automated web accessibility testing software for managers Lazar et al. International Conference on Human-Computer Interaction 2022
266 Gamified exploratory GUI testing of web applications: A preliminary evaluation Fulcini et al. Proceedings of the 2022 IEEE International Conference on Software Testing, Verification and Validation Workshops (ICSTW’22) 2022
267 Generating test paths to detect XSS vulnerabilities of web applications Nguyen et al. Proceedings of the 9th NAFOSTED Conference on Information and Computer Science (NICS’22) 2022
268 LogBERT-BiLSTM: Detecting malicious web requests Ramos Júnior et al. International Conference on Artificial Neural Networks 2022
269 Measuring consistency metric for web applications Dane et al. Proceedings of the 2022 IEEE 12th Annual Computing and Communication Workshop and Conference (CCWC’22) 2022
270 Morpheus web testing: A tool for generating test cases for widget based web applications de Almeida Neves et al. Journal of Web Engineering 2022
271 Multiperspective web testing supported by a generation hyper-heuristic Balera et al. Proceedings of the 22nd International Conference on Computational Science and Its Applications (ICCSA’22) 2022
272 Penetration testing on web application using insecure direct object references (IDOR) method Pratama et al. Proceedings of the 2022 International Conference on ICT for Smart Society (ICISS’22) 2022
273 PHP web shell detection through static analysis of AST using LSTM based deep learning Gogoi et al. Proceedings of the 1st International Conference on Artificial Intelligence Trends and Pattern Recognition (ICAITPR’22) 2022
274 Prioritising test scripts for the testing of memory bloat in web applications Qian et al. IET Software 2022
275 Semantic web Racer: Dynamic security testing of the web application against race condition in the business layer Alidoosti et al. Expert Systems with Applications 2022
276 Test-driven feature extraction of web components Long et al. Journal of Computer Science and Technology 2022
277 Web browser extension development of structured query language injection vulnerability detection using long short-term memory algorithm Turnip et al. Proceedings of the 2022 IEEE International Conference of Computer Science and Information Technology (ICOSNIKOM’22) 2022
278 Web element identification by combining NLP and heuristic search for web testing Kirinuki et al. Proceedings of the 2022 IEEE International Conference on Software Analysis, Evolution and Reengineering (SANER’22) 2022
279 Zokfuzz: Detection of web vulnerabilities via fuzzing Zhang et al. Proceedings of the 2nd International Conference on Consumer Electronics and Computer Engineering (ICCECE’22) 2022
280 A comprehensive evaluation of Q-learning based automatic web GUI testing Fan et al. Proceedings of the 10th International Conference on Dependable Systems and Their Applications (DSA’23) 2023
281 A large-scale web accessibility analysis considering technology adoption Martins et al. Universal Access in the Information Society 2023
282 A reinforcement learning approach to generating test cases for web applications Chang et al. Proceedings of the 2023 IEEE/ACM International Conference on Automation of Software Test (AST’23) 2023
283 ACUM: An approach to combining unsupervised methods for detecting malicious web sessions Demirel et al. Proceedings of the 8th International Conference on Computer Science and Engineering (UBMK’23) 2023
284 An empirical study on web GUI load testing and the hybrid with HTTP requests Lo et al. Proceedings of the 2023 IEEE International Conference on e-Business Engineering (ICEBE’23) 2023
285 An ensemble approach for visual testing of web applications Aditya et al. Proceedings of the 3rd International Conference on Emerging Frontiers in Electrical and Electronic Technologies (ICEFEET’23) 2023
286 Analytical evaluation of web performance testing tools: Apache JMeter and SoapUI Tiwari et al. Proceedings of the IEEE 12th International Conference on Communication Systems and Network Technologies (CSNT’23) 2023
287 Cytestion: Automated GUI testing for web applications Moura et al. Proceedings of the XXXVII Brazilian Symposium on Software Engineering (SBSI’23) 2023
288 DaNuoYi: Evolutionary multi-task injection testing on web application firewalls Li et al. IEEE Transactions on Software Engineering 2023
289 Deep learning-based detection of CSRF vulnerabilities in web applications Beltrano et al. Proceedings of the 2023 IEEE International Conference on Dependable, Autonomic and Secure Computing, International Conference on Pervasive Intelligence and Computing, International Conference on Cloud and Big Data Computing, International Conference on Cyber Science and Technology Congress (DASC/PiCom/CBDCom/CyberSciTech’23) 2023
290 Detecting web application DAST attacks with machine learning Shahrivar et al. Proceedings of the 2023 IEEE Conference on Dependable and Secure Computing (DSC’23) 2023
291 Detecting web-based attacks: A comparative analysis of machine learning and BERT transformer approaches Jinad et al. Proceedings of the 36th International Conference on Industrial, Engineering and Other Applications of Applied Intelligent Systems (IEA/AIE’23) 2023
292 Detection and classification of web application attacks Ramamoorthy et al. Proceedings of the 36th International Conference on Industrial, Engineering and Other Applications of Applied Intelligent Systems (IEA/AIE’23) 2023
293 E2E-Loader: A framework to support performance testing of web applications Battista et al. Proceedings of the 16th IEEE Conference on Software Testing, Verification and Validation (ICST’23) 2023
294 Effective web application vulnerability testing system using proposed XSS_SQL_Scanning_Algorithm Aung et al. Proceedings of the 2023 IEEE Conference on Computer Applications (ICCA’23) 2023
295 Efficient web application mapping during the reconnaissance phase of penetration testing Lazarov et al. Proceedings of the 15th International Congress on Ultra Modern Telecommunications and Control Systems and Workshops (ICUMT’23) 2023
296 Enhancing web application penetration testing with a static application security testing (SAST) tool Darus et al. Proceedings of the IEEE 8th International Conference on Recent Advances and Innovations in Engineering (ICRAIE’23) 2023
297 Evaluation of machine learning methods for impostor detection in web applications Grzenda et al. Expert Systems with Applications 2023
298 Facilitating web vulnerability detection on embedded devices with root path pruning Li et al. Proceedings of the 3rd International Conference on Computer Science, Electronic Information Engineering and Intelligent Control Technology (CEI’23) 2023
299 From user stories to end-to-end web testing Mollah et al. Proceedings of the 16th IEEE International Conference on Software Testing, Verification and Validation Workshops (ICSTW’23) 2023
300 Generative neural networks as a tool for web applications penetration testing Gallus et al. Proceedings of the 2023 Communication and Information Technologies (KIT’23) 2023
301 Implementation of distributed attack penetration testing automation using dynamic infrastructure framework Axiom on web-based systems Prisadi et al. Proceedings of the 2023 International Conference on Informatics, Multimedia, Cyber and Informations System (ICIMCIS’23) 2023
302 Machine learning based web application plugin for threat detection and IP analysis Katulanda et al. Proceedings of the 2023 5th International Conference on Advancements in Computing (ICAC’23) 2023
303 Manual and automated web-based testing of the NPTEL portal DA et al. Proceedings of the 2023 International Conference on Artificial Intelligence and Knowledge Discovery in Concurrent Engineering (ICECONF’23) 2023
304 Metamorphic testing for web system security Chaleshtari et al. IEEE Transactions on Software Engineering 2023
305 Parallel evolutionary test case generation for web applications Wang et al. Information and Software Technology 2023
306 RanABD: MTD-based technique for detection of advanced session-replay web bots Sadeghpour et al. Proceedings of the 10th ACM Workshop on Moving Target Defense (MTD’23) 2023
307 Scanner++: Enhanced vulnerability detection of web applications with attack intent synchronization Yin et al. ACM Transactions on Software Engineering and Methodology 2023
308 Splendor: static detection of stored XSS in modern web applications Su et al. Proceedings of the 32nd ACM SIGSOFT International Symposium on Software Testing and Analysis (ISSTA’23 2023
309 Web application penetration testing through rule-based inference on knowledge graphs Caturano et al. Proceedings of the 2023 International Conference on Electrical, Computer and Energy Technologies (ICECET’23) 2023
310 Web application security testing framework using flask Rohit et al. Proceedings of the 2nd International Conference on Applied Artificial Intelligence and Computing (ICAAIC’23) 2023
311 WebEV: A dataset on the behavior of testers for web application end to end testing Fuad et al. Proceedings of the IEEE/ACM 31st International Conference on Program Comprehension (ICPC’23) 2023
312 Artificial intelligence web application firewall for advanced detection of web injection attacks Román-Gallego et al. Expert Systems 2023
313 Generative pre-trained transformer-based reinforcement learning for testing web application firewalls Liang et al. IEEE Transactions on Dependable and Secure Computing 2023
314 Mutta: A novel tool for E2E web mutation testing Leotta et al. Software Quality Journal 2023